### v1

# Stable Distributed Online Feedback Optimization for Distribution System Voltage Regulation

Abstract
We investigate the distributed voltage regulation problem in distribution systems employing online feedback optimization and short-range communication between physical neighbours. We show that a two-metric approach can be unstable. As a remedy, we propose a nested feedback optimization strategy. Simulation results reveal that while the two-metric approach fails to regulate voltages, the proposed approach achieves even less voltage limit violations than its centralized counterpart.
Index Terms:
Distributed communication, online feedback optimization, voltage regulation, reactive power control
I Introduction
With the proliferation of distributed energy resources (DERs) in distribution systems, voltage regulation has become a challenging task for distribution system operators (DSOs). While local control cannot guarantee the desired voltage regulation [1], model-based feedforward optimization requires accurate grid models, real-time load measurements, and lacks robustness to system uncertainties and disturbances [2].
Recently, online feedback optimization (OFO) has emerged as a promising approach for real-time distribution system control. OFO utilizes measurements as feedback and employs optimization algorithms as feedback controllers to steer the system towards an optimal operating point [2]. While many OFO controllers rely on centralized communication [3, 4], distributed OFO, which relies on communication between physical neighbours, appears to be a compelling approach due to its robustness to single-point failures [5, 1, 6, 7, 8].
While [5, 1, 6] demand particular objective function designs, [7] relies on disseminating global information through neighboring nodes. However, the systematic delays inherent in this information propagation may raise stability issues, potentially undermining its efficacy in practical distribution systems. In this letter, we propose a distributed OFO design that is closely related to that in [8], where the distributed communication is achieved by scaling the gradient with a sparse positive definite matrix $\mathbf{X}^{-1}$. A two-metric approach [9] was further used in [8], where an Euclidean projection step was used to ensure satisfaction of the DER capacity constraint. The approach is called two-metric because the matrix $\mathbf{X}^{-1}$ is used for the gradient scaling while the identity matrix $\mathbf{I}$ is used for the projection.
However, utilizing such a two-metric approach might lead to divergence, as shown in [9]. Instead of the Euclidean distance, the distance measured in terms of the norm $\lVert\mathbf{z}\rVert_{\mathbf{X}}=\sqrt{\mathbf{z}^{\intercal}\mathbf{X}\mathbf{z}}$ should be used in the projection step to ensure convergence [9]. Solving this non-Euclidean projection problem would however require global information, contradicting the distributed OFO design. To tackle this challenge, we propose an iterative algorithm that requires only local information and operates in an online feedback-based manner. This leads to our nested feedback optimization approach, where the outer loop comprises the OFO iterations yielding tentative but not necessarily feasible DER setpoints, and the inner loop solves the non-Euclidean projection to map the tentative setpoints to actual feasible setpoints.
II Online feedback optimization
II-A Problem formulation
Consider the distribution system voltage regulation using reactive power as in Problem (1). DERs are located in the set $\mathcal{N}$, and $q_{i}$ represents their reactive power injections. The objective function (1a) minimizes the reactive power cost, where $c_{i}$ is the cost parameter. Constraint (1b) enforces the lower and upper voltage limits $\underline{v}$ and $\overline{v}$, where $\mu_{i}$ and $\lambda_{i}$ are the respective dual variables. Finally, (1c) ensures that the reactive power injection $q_{i}$ respects the lower and upper limits $\underline{q}_{i}$ and $\overline{q}_{i}$. Note that the voltage $v_{i}$ is dependent on $q_{i}$ and other disturbances in the system such as active power injections of the DERs and active and reactive power consumption of non-controllable load units.
$\underset{q_{i},\forall i\in\mathcal{N}}{\operatorname{minimize}}$ $\quad f:=\frac{1}{2}c_{i}q_{i}^{2}$ (1a)
$\operatorname{s.t.}$ $\quad\underline{v}\leq v_{i}\leq\overline{v}:\mu_{i},\lambda_{i},\forall i\in\mathcal{N}$ (1b)
$\underline{q}_{i}\leq q_{i}\leq\overline{q}_{i},\forall i\in\mathcal{N}$ (1c)
II-B Centralized primal-dual gradient projection algorithm
Solving Problem (1) offline in a feedforward way demands an accurate grid model and disturbance data to evaluate $v_{i}$, which might not be available. The feedforward approach also lacks robustness [2]. To tackle these challenges, OFO has emerged as a promising approach for real-time distribution system management. This section introduces an instance of OFO based on the primal-dual gradient projection (PDGP) algorithm [3], which requires a centralized gather-and-broadcast communication architecture. The PDGP algorithm performs projected gradient ascent and descent steps iteratively for the dual and primal variables, respectively. At each time step (iteration) $k$, it repeats the following three steps:
  1. 1.
    For each bus $i\in\mathcal{N}$, collect its voltage measurement $\tilde{v}_{i}^{k}$ and update $\lambda_{i}$ and $\mu_{i}$, where the projection operator is defined as $[u]^{+}:=\max(u,0)$ and $\alpha$ with various superscripts represents step sizes:
    $\lambda_{i}^{k+1}\leftarrow\big{[}\lambda_{i}^{k}+\alpha^{\lambda}(\tilde{v}_{i}^{k}-\overline{v})\big{]}^{+}$ (2a)
    $\mu_{i}^{k+1}\leftarrow\big{[}\mu_{i}^{k}+\alpha^{\mu}(\underline{v}-\tilde{v}_{i}^{k})\big{]}^{+}$ (2b)
  2. 2.
    The DSO broadcasts the updated dual variables $\boldsymbol{\lambda}^{k+1}$ and $\boldsymbol{\mu}^{k+1}$ to all buses, where $\boldsymbol{\lambda}^{k+1}$ and $\boldsymbol{\mu}^{k+1}$ are vectors constructed by stacking $\lambda_{i}^{k+1}$ and $\mu_{i}^{k+1}$, respectively.
  3. 3.
    For each bus $i\in\mathcal{N}$, update its reactive power setpoints $q_{i}$, where $[u]_{\underline{q}_{i}}^{\overline{q}_{i}}$ represents the Euclidean projection into $[\underline{q}_{i},\overline{q}_{i}]$, and $\mathbf{X}$ is a $|\mathcal{N}|$-dimensional dense matrix mapping reactive power injections to voltages, i.e., the sensitivity matrix, as constructed in [8]. Because of the non-sparsity of $\mathbf{X}$, Step 2 is required to provide information needed in Step 3.
    $q_{i}^{k+1}\leftarrow\big{[}q_{i}^{k}-\alpha[c_{i}q_{i}^{k}+[\mathbf{X}(\boldsymbol{\lambda}^{k+1}-\boldsymbol{\mu}^{k+1})]_{i}]\big{]}_{\underline{q}_{i}^{k}}^{\overline{q}_{i}^{k}}$ (2c)
II-C Two-metric distributed approach
To avoid a centralized communication architecture and promote the distributed communication architecture where communication is only between physical neighbours in the system, [8] proposed to scale the gradient in (2c) by a sparse positive definite matrix $\mathbf{X}^{-1}$, which has non-zero entries only in the diagonal and the $i$-th row $j$-th column if nodes $i$ and $j$ are adjacent. This yields (3) and replaces (2c) in the OFO, where $\mathbf{q}^{k}$ is the stack of $q^{k}_{i}$, and $\mathbf{C}$ is a $|\mathcal{N}|$-dimensional diagonal matrix with elements $c_{i}$ defined in (1a).
$q_{i}^{k+1}\leftarrow\big{[}q_{i}^{k}-\alpha[[\mathbf{X}^{-1}\mathbf{C}\mathbf{q}^{k}]_{i}+\lambda_{i}^{k+1}-\mu_{i}^{k+1}]\big{]}_{\underline{q}_{i}^{k}}^{\overline{q}_{i}^{k}}$ (3)
Computing (3) for node $i$ now requires only information from its physcial neighbours. We stack $q_{i}^{k}-\alpha[\mathbf{X}^{-1}\mathbf{C}\mathbf{q}^{k}]_{i}+\lambda_{i}^{k+1}-\mu_{i}^{k+1}$ for all nodes (tentative but not necessarily feasible setpoints to be projected) as the vector $\mathbf{\dot{q}}^{k+1}$ and define $\mathcal{X}^{k}$ as $[\underline{q}_{1}^{k},\overline{q}_{1}^{k}]\times[\underline{q}_{2}^{k},\overline{q}_{2}^{k}]\times\cdots\times[\underline{q}_{|\mathcal{N}|}^{k},\overline{q}_{|\mathcal{N}|}^{k}]$ for further developments, where $\times$ is the Cartesian product of sets.
II-D Proposed distributed approach
The two-metric approach in Section II-C does not necessarily lead to algorithm convergence [9]. To ensure convergence, the projection in (3) should use the distance measured in terms of the norm $\lVert\mathbf{z}\rVert_{\mathbf{X}}=\sqrt{\mathbf{z}^{\intercal}\mathbf{X}\mathbf{z}}$ due to the use of the scaling matrix $\mathbf{X}^{-1}$ [9], that is:
$\mathbf{q}^{k+1}\leftarrow\underset{\mathbf{u}\in\mathcal{X}^{k}}{\operatorname{argmin}}\quad\frac{1}{2}(\mathbf{u}-\mathbf{\dot{q}}^{k+1})^{\intercal}\mathbf{X}(\mathbf{u}-\mathbf{\dot{q}}^{k+1})$ (4)
This however contradicts the distributed communication architecture of the overall OFO algorithm since directly solving (4) again requires global information. To tackle this challenge, this letter proposes an iterative algorithm that requires only local information and is implemented in an online feedback-based manner. This leads to our nested feedback optimization approach, where the outer loop comprises the OFO iterations giving tentative reactive power setpoints, and the inner loop solves the non-Euclidean projection problem (4). Our key observation is that given a feasible $\mathbf{u}^{\tau}$, the gradient of the objective function in (4) is $\mathbf{X}(\mathbf{u}^{\tau}-\mathbf{\dot{q}}^{k+1})$, which can be approximated by the voltage difference between $\mathbf{v}(\mathbf{u}^{\tau})$ and $\mathbf{v}(\mathbf{\dot{q}}^{k+1})$ assuming system disturbances do not change significantly when all nodes are locally implementing this inner iterative algorithm. The algorithm is based on the non-scaled Euclidean gradient projection and proceeds as:
  1. 1.
    Implement $\mathbf{q}^{k}_{\epsilon}:=\mathbf{q}^{k}+\epsilon(\mathbf{\dot{q}}^{k+1}-\mathbf{q}^{k})$ and measure the voltage $\mathbf{\tilde{v}}(\mathbf{q}^{k}_{\epsilon})$. Estimate $\mathbf{\tilde{v}}(\mathbf{\dot{q}}^{k+1})$ as $\mathbf{\tilde{v}}^{k}$ + ($\mathbf{\tilde{v}}(\mathbf{q}^{k}_{\epsilon})$-$\mathbf{\tilde{v}}^{k}$)/$\epsilon$.
  2. 2.
    Given a feasible $\mathbf{u}^{\tau}$, implement it, measure the voltage $\mathbf{\tilde{v}}(\mathbf{u}^{\tau})$, compute the next implementable reactive power setpoint $\mathbf{u}^{\tau+1}$ as (5), and repeat this step for $T$ iterations.
    $\mathbf{u}^{\tau+1}\leftarrow[\mathbf{u}^{\tau}-\alpha^{u}(\mathbf{\tilde{v}}(\mathbf{u}^{\tau})-\mathbf{\tilde{v}}(\mathbf{\dot{q}}^{k+1}))]_{\mathbf{\underline{q}}^{k}}^{\mathbf{\overline{q}}^{k}}$ (5)
Note that these two steps can be implemented for each node using only local information. In the first step, $\epsilon$ represents a small exploration parameter, which is similar to that in model-free feedback optimization [10]. Since $\mathbf{\dot{q}}^{k+1}$ can be pretty far away from the feasible set, it might be impossible to implement it. Therefore, we implement $\mathbf{q}^{k}_{\epsilon}$ and estimate $\mathbf{\tilde{v}}(\mathbf{\dot{q}}^{k+1})$ based on it. This could still result in a minor temporary capacity limit violation, which should be manageable by the DER. To ensure strict capacity limit satisfaction, the original feasible set can be slightly deflated [10]. The initial feasible $\mathbf{u}^{\tau}$ in the second step can be simply chosen as $\mathbf{q}^{k}$.
III Simulation results
In our simulation experiment, we compare the centralized OFO based on the gather-and-broadcast communication, the distributed two-metric approach in [8], and our proposed nested distributed approach in both static and dynamic cases. We study a 100% penetration scenario of photovoltaics (PVs) in a 96-node low-voltage distribution system, which experiences overvoltage issues. Details of the distribution system and the generation and load data with a 6-second resolution can be found in [11] and references therein. The PV inverters are assumed to be oversized by 20% to provide sufficient reactive power. The reactive power cost matrix $\mathbf{C}$ is chosen as the identity matrix. The lower and upper voltage limits are 0.95 and 1.05 pu, respectively. The unit for power is $kW/kVar$. The step sizes are chosen with a trial-and-error strategy [8] as $\alpha^{\lambda}=\alpha^{\mu}=10^{6}$, $\alpha=10^{-1}$ (for centralized), $\alpha=10^{-3}$ (for distributed), and $\alpha^{u}=10^{2}$. The exploration parameter $\epsilon$ is chosen as $10^{-5}$, resulting in a maximum DER reactive power limit violation of 2%. For each generation and load data point, an OFO iteration is run which yields tentative reactive power setpoints, followed by one exploration step and 5 inner feedback iterations of (5) ($T=5$).
Figure 1: Voltage profiles over 200 iterations for two remote buses in the static case using the proposed and the two-metric approaches. We count both the outer and inner iterations for our proposed approach.
Figure 2: Reactive power profiles after 200 iterations for the centralized, the proposed, and the two-metric approaches.
Figure 3: Voltage profiles for the most remote bus in the dynamic simulation. The average voltage violation (AVV) values are $1.9\times 10^{-4}$, $1.3\times 10^{-4}$, and $4.6\times 10^{-3}$ pu for the centralized, the proposed, and the two-metric approaches, respectively, which is defined as $\operatorname{AVV}=\frac{1}{K}\sum_{k=1}^{K}[v^{k}-\overline{v}]^{+}$, where $k$ is the iteration index and $K$ is the number of iterations.
In the static case, the generation and load data are kept unchanged to study the convergence of various approaches. Figure 1 shows that while the two-metric approach does not converge and fails to regulate voltages, our proposed approach converges quickly after approximately 154 iterations, which correspond to around 22 outer and 132 inner iterations. We also compare the final reached reactive power setpoints after 200 iterations. Figure 2 shows that the centralized and our proposed approaches converge to the same solution. This is reasonable since the gradient scaling only alters the trajectory, but does not change the solution to Problem (1).
In the dynamic case, we study the performance of the approaches in a time-varying environment, where the approaches are implemented online without reaching convergence. Figure 3 shows the significant voltage limit violation without control and demonstrates that our proposed nested distributed approach successfully enforces the voltage limit even for the most remote, and therefore, most sensitive bus, achieving an average voltage violation (AVV) of $1.3\times 10^{-4}$ pu, which is better even than that of the centralized approach ($1.9\times 10^{-4}$ pu). We conjecture that this is because of the smoothing brought by the inner feedback loop. The two-metric approach, however, fails to regulate voltages, resulting in 34 times higher AVV.
IV Conclusion
In this letter, we focused on the distributed voltage regulation problem in distribution systems. We showed that the previously proposed two-metric approach might fail. As a remedy, we proposed a nested feedback optimization approach. Our simulation results showed that the approach achieved even better voltage regulation than a centralized approach while only requiring short-range communication between physical neighbours. The proposed distributed approach however does not extend to congestion management while the centralized does so. The approach works for either active or reactive power control, but not joint control. This is because the observation in Section II-D does not hold anymore since the voltage difference is no longer the gradient of either active or reactive power injection, but represents their joint non-separable impact. However, one might deploy active and reactive power feedback controllers separately at different sampling time. Whether the system under two feedback controllers is stable is left for future investigations.

### v2

@@
-Stable Distributed Online Feedback Optimization for Distribution System Voltage Regulation
+Distributed Online Feedback Optimization for Real-time Distribution System Voltage Regulation
 Abstract
-We investigate the distributed voltage regulation problem in distribution systems employing online feedback optimization and short-range communication between physical neighbours.
-We show that a two-metric approach can be unstable.
-As a remedy, we propose a nested feedback optimization strategy.
-Simulation results reveal that while the two-metric approach fails to regulate voltages, the proposed approach achieves even less voltage limit violations than its centralized counterpart.
+We investigate the real-time distribution system voltage regulation problem employing online feedback optimization (OFO) and short-range communication between physical neighbours.
+OFO does not need an accurate grid topology nor estimated consumption of non-controllable loads, affords fast calculations, and demonstrates robustness to uncertainties due to its feedback-based nature.
+These characteristics render it particularly suitable for real-time distribution system applications.
+However, many OFO controllers require centralized communication, which lack robustness to single-point failures.
+This paper proposes a distributed OFO design based on a nested feedback optimization strategy and analyzes its convergence.
+Numerical results reveal that the proposed design achieves satisfactory voltage regulation and outperforms other distributed approaches.
 Index Terms:
@@
 I Introduction
-With the proliferation of distributed energy resources (DERs) in distribution systems, voltage regulation has become a challenging task for distribution system operators (DSOs).
-While local control cannot guarantee the desired voltage regulation [1], model-based feedforward optimization requires accurate grid models, real-time load measurements, and lacks robustness to system uncertainties and disturbances [2].
-Recently, online feedback optimization (OFO) has emerged as a promising approach for real-time distribution system control.
-OFO utilizes measurements as feedback and employs optimization algorithms as feedback controllers to steer the system towards an optimal operating point [2].
-While many OFO controllers rely on centralized communication [3, 4], distributed OFO, which relies on communication between physical neighbours, appears to be a compelling approach due to its robustness to single-point failures [5, 1, 6, 7, 8].
-While [5, 1, 6] demand particular objective function designs, [7] relies on disseminating global information through neighboring nodes.
+With the proliferation of distributed energy resources (DERs) in distribution systems, real-time voltage regulation has become a challenging task for distribution system operators (DSOs).
+While local control cannot guarantee the desired voltage regulation [1], model-based feedforward optimization requires an accurate grid model, load measurements, and lacks robustness to system uncertainties and disturbances [2].
+To overcome these limitations, online feedback optimization (OFO) has been proposed as a promising solution approach.
+OFO utilizes measurements as feedback and employs optimization algorithms as feedback controllers to steer physical systems towards their optimal operating points [2]; see Fig.
+1 for comparison between feedforward and feedback systems.
+The feedback-based implementation eliminates the need for an accurate grid model and real-time load data and brings robustness to OFO.
+Compared to deep reinforcement learning, OFO does not need a complicated training process and has guarantees to pursue grid-level objectives.
+Compared to using traditional voltage regulators such as on-load tap changers and capacitor banks, OFO provides fast voltage regulation and can complement these slow-reacting devices in real time [3].
+While many OFO controllers rely on centralized communication [4, 5, 6, 7, 8], distributed OFO based on communication between physical neighbours appears a compelling approach due to its robustness to single-point failures[9, 1, 3, 10, 11, 12].
+Among the existing distributed designs, [9, 1, 3] demand particular forms of objective functions, while [10] relies on disseminating global information through neighboring nodes.
 However, the systematic delays inherent in this information propagation may raise stability issues, potentially undermining its efficacy in practical distribution systems.
-In this letter, we propose a distributed OFO design that is closely related to that in [8], where the distributed communication is achieved by scaling the gradient with a sparse positive definite matrix $\mathbf{X}^{-1}$.
-A two-metric approach [9] was further used in [8], where an Euclidean projection step was used to ensure satisfaction of the DER capacity constraint.
-The approach is called two-metric because the matrix $\mathbf{X}^{-1}$ is used for the gradient scaling while the identity matrix $\mathbf{I}$ is used for the projection.
-However, utilizing such a two-metric approach might lead to divergence, as shown in [9].
-Instead of the Euclidean distance, the distance measured in terms of the norm $\lVert\mathbf{z}\rVert_{\mathbf{X}}=\sqrt{\mathbf{z}^{\intercal}\mathbf{X}\mathbf{z}}$ should be used in the projection step to ensure convergence [9].
-Solving this non-Euclidean projection problem would however require global information, contradicting the distributed OFO design.
-To tackle this challenge, we propose an iterative algorithm that requires only local information and operates in an online feedback-based manner.
-This leads to our nested feedback optimization approach, where the outer loop comprises the OFO iterations yielding tentative but not necessarily feasible DER setpoints, and the inner loop solves the non-Euclidean projection to map the tentative setpoints to actual feasible setpoints.
+In [11], the distributed design was achieved by only keeping components in the network sensitivity matrix related to local and neighbouring nodes.
+This however does not assure voltage regulation.
+In this paper, we propose a distributed OFO design related to that in [12], where the distributed implementation was enabled by the sparsity of the inverse of the network sensitivity matrix $\mathbf{X}$, i.e.
+$\mathbf{X}^{-1}$ is sparse.
+Both approaches continue with scaling the gradient by $\mathbf{X}^{-1}$ followed by a projection step to ensure satisfaction of the DER capacity constraint.
+It is conceptually attractive to use the Euclidean distance in the projection, which can be implemented using only local information.
+However, as shown in [13], the distance measured by the norm $\lVert\mathbf{z}\rVert_{\mathbf{X}}=\sqrt{\mathbf{z}^{\intercal}\mathbf{X}\mathbf{z}}$ should be used instead in the projection step to ensure descent iterations and thus algorithm convergence.
+With the Euclidean distance, it would be a two-metric approach which can lead to algorithm divergence [13].
+Solving this non-Euclidean projection problem would however require global information since $\mathbf{X}$ is dense, contradicting the distributed OFO design.
+As a remedy, [12] introduced additional dualization of the DER capacity limits on top of the Euclidean projection to ensure the DER constraint satisfaction.
+Shown in their numerical result and our simulation, this strategy takes a prohibitively large number of iterations to converge, which jeopardizes its performance in online implementation.
+Figure 1: Block diagrams of feedforward and feedback systems for voltage regulation using reactive power, where $q$ represents system input (reactive power), $v$ represents system output (voltages), $\omega$ ($\widehat{\omega}$) represents (estimated) system disturbance (non-controllable loads), and $h$ maps system input and disturbance to system output.
+We tackle this challenge by proposing an iterative algorithm to solve this non-Euclidean projection problem that requires only local information and operates in an online feedback-based manner.
+This leads to our overall nested online feedback optimization approach, where the outer loop comprises the OFO iterations requiring communication between physical neighbours and yielding tentative but not necessarily feasible DER setpoints, and the inner loop solves the non-Euclidean projection problem to map the tentative setpoints to actual feasible setpoints using local information.
+The remainder of this paper is structured as follows: Section II presents the problem formulation and the proposed nested approach and analyzes its convergence.
+Section III presents simulation results, while Section IV draws conclusions and discusses future work.
 II Online feedback optimization
 II-A Problem formulation
-Consider the distribution system voltage regulation using reactive power as in Problem (1).
-DERs are located in the set $\mathcal{N}$, and $q_{i}$ represents their reactive power injections.
-The objective function (1a) minimizes the reactive power cost, where $c_{i}$ is the cost parameter.
+Consider the distribution system voltage regulation problem using reactive power as formulated in (1).
+DERs are located in the set $\mathcal{N}$, and $q_{i}$ represents their respective reactive power injection.
+The objective function (1a) minimizes the reactive power use cost, where $c_{i}>0$ is the cost parameter.
 Constraint (1b) enforces the lower and upper voltage limits $\underline{v}$ and $\overline{v}$, where $\mu_{i}$ and $\lambda_{i}$ are the respective dual variables.
 Finally, (1c) ensures that the reactive power injection $q_{i}$ respects the lower and upper limits $\underline{q}_{i}$ and $\overline{q}_{i}$.
-Note that the voltage $v_{i}$ is dependent on $q_{i}$ and other disturbances in the system such as active power injections of the DERs and active and reactive power consumption of non-controllable load units.
+Note that $v_{i}$ is dependent on $q_{i}$ and other disturbances in the system such as active and reactive power consumption of non-controllable load units.
 $\underset{q_{i},\forall i\in\mathcal{N}}{\operatorname{minimize}}$ $\quad f:=\frac{1}{2}c_{i}q_{i}^{2}$ (1a)
@@
 II-B Centralized primal-dual gradient projection algorithm
-Solving Problem (1) offline in a feedforward way demands an accurate grid model and disturbance data to evaluate $v_{i}$, which might not be available.
+Solving (1) offline in a feedforward way demands an accurate grid model and disturbance data to evaluate $v_{i}$, which might not be available.
 The feedforward approach also lacks robustness [2].
-To tackle these challenges, OFO has emerged as a promising approach for real-time distribution system management.
-This section introduces an instance of OFO based on the primal-dual gradient projection (PDGP) algorithm [3], which requires a centralized gather-and-broadcast communication architecture.
+Thus, this section introduces an instance of OFO based on the primal-dual gradient projection (PDGP) algorithm [4, 5] to track the time-varying optimizer of (1), which requires a centralized gather-and-broadcast communication architecture.
 The PDGP algorithm performs projected gradient ascent and descent steps iteratively for the dual and primal variables, respectively.
@@
   1. 1.
-    For each bus $i\in\mathcal{N}$, collect its voltage measurement $\tilde{v}_{i}^{k}$ and update $\lambda_{i}$ and $\mu_{i}$, where the projection operator is defined as $[u]^{+}:=\max(u,0)$ and $\alpha$ with various superscripts represents step sizes:
-    $\lambda_{i}^{k+1}\leftarrow\big{[}\lambda_{i}^{k}+\alpha^{\lambda}(\tilde{v}_{i}^{k}-\overline{v})\big{]}^{+}$ (2a)
-    $\mu_{i}^{k+1}\leftarrow\big{[}\mu_{i}^{k}+\alpha^{\mu}(\underline{v}-\tilde{v}_{i}^{k})\big{]}^{+}$ (2b)
+    For each bus $i\in\mathcal{N}$, collect its voltage measurement $\tilde{v}_{i}^{k}$ and update $\lambda_{i}$ and $\mu_{i}$, where the projection operator is defined as $[u]^{+}:=\max(u,0)$, $\alpha$ with various superscripts represents step sizes, while $r^{d}$ and $r^{p}$ in (2c) are the dual and primal regularization factors [4, 5], respectively.
+    $\lambda_{i}^{k+1}\leftarrow\big{[}\lambda_{i}^{k}+\alpha^{d}(\tilde{v}_{i}^{k}-\overline{v}-r^{d}\lambda_{i}^{k})\big{]}^{+}$ (2a)
+    $\mu_{i}^{k+1}\leftarrow\big{[}\mu_{i}^{k}+\alpha^{d}(\underline{v}-\tilde{v}_{i}^{k}-r^{d}\mu_{i}^{k})\big{]}^{+}$ (2b)
   2. 2.
@@
   3. 3.
-    For each bus $i\in\mathcal{N}$, update its reactive power setpoints $q_{i}$, where $[u]_{\underline{q}_{i}}^{\overline{q}_{i}}$ represents the Euclidean projection into $[\underline{q}_{i},\overline{q}_{i}]$, and $\mathbf{X}$ is a $|\mathcal{N}|$-dimensional dense matrix mapping reactive power injections to voltages, i.e., the sensitivity matrix, as constructed in [8].
-Because of the non-sparsity of $\mathbf{X}$, Step 2 is required to provide information needed in Step 3.
-    $q_{i}^{k+1}\leftarrow\big{[}q_{i}^{k}-\alpha[c_{i}q_{i}^{k}+[\mathbf{X}(\boldsymbol{\lambda}^{k+1}-\boldsymbol{\mu}^{k+1})]_{i}]\big{]}_{\underline{q}_{i}^{k}}^{\overline{q}_{i}^{k}}$ (2c)
+    For each bus $i\in\mathcal{N}$, update its reactive power setpoints $q_{i}$, where $[u]_{\underline{q}_{i}}^{\overline{q}_{i}}$ represents the Euclidean projection into $[\underline{q}_{i},\overline{q}_{i}]$, $\mathbf{X}$ is a $|\mathcal{N}|$-dimensional dense matrix mapping reactive power injections to voltages, i.e.
+the sensitivity matrix, as constructed in [12], and $\mathbf{q}^{k}$ is the stack of $q^{k}_{i}$.
+Since $\mathbf{X}$ is dense, Step 2 is needed to provide information for Step 3.
+    $q_{i}^{k+1}\leftarrow\big{[}q_{i}^{k}-\alpha[c_{i}q_{i}^{k}+[\mathbf{X}(\boldsymbol{\lambda}^{k+1}-\boldsymbol{\mu}^{k+1}+r^{p}\mathbf{q}^{k})]_{i}]\big{]}_{\underline{q}_{i}^{k}}^{\overline{q}_{i}^{k}}$ (2c)
 II-C Two-metric distributed approach
-To avoid a centralized communication architecture and promote the distributed communication architecture where communication is only between physical neighbours in the system, [8] proposed to scale the gradient in (2c) by a sparse positive definite matrix $\mathbf{X}^{-1}$, which has non-zero entries only in the diagonal and the $i$-th row $j$-th column if nodes $i$ and $j$ are adjacent.
-This yields (3) and replaces (2c) in the OFO, where $\mathbf{q}^{k}$ is the stack of $q^{k}_{i}$, and $\mathbf{C}$ is a $|\mathcal{N}|$-dimensional diagonal matrix with elements $c_{i}$ defined in (1a).
-$q_{i}^{k+1}\leftarrow\big{[}q_{i}^{k}-\alpha[[\mathbf{X}^{-1}\mathbf{C}\mathbf{q}^{k}]_{i}+\lambda_{i}^{k+1}-\mu_{i}^{k+1}]\big{]}_{\underline{q}_{i}^{k}}^{\overline{q}_{i}^{k}}$ (3)
-Computing (3) for node $i$ now requires only information from its physcial neighbours.
-We stack $q_{i}^{k}-\alpha[\mathbf{X}^{-1}\mathbf{C}\mathbf{q}^{k}]_{i}+\lambda_{i}^{k+1}-\mu_{i}^{k+1}$ for all nodes (tentative but not necessarily feasible setpoints to be projected) as the vector $\mathbf{\dot{q}}^{k+1}$ and define $\mathcal{X}^{k}$ as $[\underline{q}_{1}^{k},\overline{q}_{1}^{k}]\times[\underline{q}_{2}^{k},\overline{q}_{2}^{k}]\times\cdots\times[\underline{q}_{|\mathcal{N}|}^{k},\overline{q}_{|\mathcal{N}|}^{k}]$ for further developments, where $\times$ is the Cartesian product of sets.
+To avoid the centralized communication requirement and enable the distributed communication architecture where communication is only between physical neighbours, [12] proposed to scale the gradient in (2c) by a sparse positive definite matrix $\mathbf{X}^{-1}$, which has non-zero entries only in the diagonal and the $i$-th row $j$-th column if buses $i$ and $j$ are adjacent.
+Continuing with the Euclidean projection yields (3) which replaces (2c) in the OFO where $\mathbf{C}$ is a $|\mathcal{N}|$-dimensional diagonal matrix with $c_{i}$ defined in (1a).
+$q_{i}^{k+1}\leftarrow\big{[}q_{i}^{k}-\alpha[[\mathbf{X}^{-1}\mathbf{C}\mathbf{q}^{k}]_{i}+\lambda_{i}^{k+1}-\mu_{i}^{k+1}+r^{p}q_{i}^{k}]\big{]}_{\underline{q}_{i}^{k}}^{\overline{q}_{i}^{k}}$ (3)
+Computing (3) for node $i$ now requires only information exchange with its physical neighbours.
+We stack $q_{i}^{k}-\alpha[[\mathbf{X}^{-1}\mathbf{C}\mathbf{q}^{k}]_{i}+\lambda_{i}^{k+1}-\mu_{i}^{k+1}+r^{p}q_{i}^{k}]$ for all nodes (tentative but not necessarily feasible setpoints to be projected) as the vector $\mathbf{\dot{q}}^{k+1}$ and define $\mathcal{X}^{k}$ as $[\underline{q}_{1}^{k},\overline{q}_{1}^{k}]\times[\underline{q}_{2}^{k},\overline{q}_{2}^{k}]\times\cdots\times[\underline{q}_{|\mathcal{N}|}^{k},\overline{q}_{|\mathcal{N}|}^{k}]$, where $\times$ is the Cartesian product of sets.
 II-D Proposed distributed approach
-The two-metric approach in Section II-C does not necessarily lead to algorithm convergence [9].
-To ensure convergence, the projection in (3) should use the distance measured in terms of the norm $\lVert\mathbf{z}\rVert_{\mathbf{X}}=\sqrt{\mathbf{z}^{\intercal}\mathbf{X}\mathbf{z}}$ due to the use of the scaling matrix $\mathbf{X}^{-1}$ [9], that is:
+The two-metric approach in Section II-C does not necessarily lead to algorithm convergence since it is not in general a descent iteration [13].
+To ensure convergence, the projection in (3) should use the distance measured in terms of the norm $\lVert\mathbf{z}\rVert_{\mathbf{X}}=\sqrt{\mathbf{z}^{\intercal}\mathbf{X}\mathbf{z}}$ due to the scaling matrix $\mathbf{X}^{-1}$ [13], that is:
 $\mathbf{q}^{k+1}\leftarrow\underset{\mathbf{u}\in\mathcal{X}^{k}}{\operatorname{argmin}}\quad\frac{1}{2}(\mathbf{u}-\mathbf{\dot{q}}^{k+1})^{\intercal}\mathbf{X}(\mathbf{u}-\mathbf{\dot{q}}^{k+1})$ (4)
-This however contradicts the distributed communication architecture of the overall OFO algorithm since directly solving (4) again requires global information.
-To tackle this challenge, this letter proposes an iterative algorithm that requires only local information and is implemented in an online feedback-based manner.
-This leads to our nested feedback optimization approach, where the outer loop comprises the OFO iterations giving tentative reactive power setpoints, and the inner loop solves the non-Euclidean projection problem (4).
-Our key observation is that given a feasible $\mathbf{u}^{\tau}$, the gradient of the objective function in (4) is $\mathbf{X}(\mathbf{u}^{\tau}-\mathbf{\dot{q}}^{k+1})$, which can be approximated by the voltage difference between $\mathbf{v}(\mathbf{u}^{\tau})$ and $\mathbf{v}(\mathbf{\dot{q}}^{k+1})$ assuming system disturbances do not change significantly when all nodes are locally implementing this inner iterative algorithm.
-The algorithm is based on the non-scaled Euclidean gradient projection and proceeds as:
+This however contradicts the distributed communication architecture of the overall OFO algorithm since, unlike the Euclidean projection (equivalent to the case where $\mathbf{X}$ is the identity matrix), solving (4) again requires global information.
+As a solution, [12] proposed dualization of the DER capacity limits, combined with the Euclidean projection, to ensure the DER constraint satisfaction.
+However, as demonstrated both in their numerical result and our simulation, this approach requires an excessively large number of iterations to converge, significantly hindering its performance in online implementation.
+To tackle this challenge, this paper proposes an iterative algorithm to solve (4) that requires only local information and is implemented in an online feedback-based manner.
+This leads to our nested feedback optimization approach, where the outer loop comprises the OFO iterations giving tentative reactive power setpoints, and the inner loop solves the non-Euclidean projection problem (4) to generate feasible setpoints.
+Our key observation is that given a feasible $\mathbf{u}^{\tau}\in\mathcal{X}^{k}$, the gradient of the objective function in (4) at $\mathbf{u}^{\tau}$ is $\mathbf{X}(\mathbf{u}^{\tau}-\mathbf{\dot{q}}^{k+1})$, which can be approximated by the voltage difference $\mathbf{v}(\mathbf{u}^{\tau})-\mathbf{v}(\mathbf{\dot{q}}^{k+1})$ assuming system disturbances do not change significantly when all nodes are locally implementing this inner iterative algorithm.
+The proposed algorithm is based on the non-scaled Euclidean gradient projection and proceeds as:
   1. 1.
@@
     $\mathbf{u}^{\tau+1}\leftarrow[\mathbf{u}^{\tau}-\alpha^{u}(\mathbf{\tilde{v}}(\mathbf{u}^{\tau})-\mathbf{\tilde{v}}(\mathbf{\dot{q}}^{k+1}))]_{\mathbf{\underline{q}}^{k}}^{\mathbf{\overline{q}}^{k}}$ (5)
-Note that these two steps can be implemented for each node using only local information.
 In the first step, $\epsilon$ represents a small exploration parameter, which is similar to that in model-free feedback optimization [14].
@@
 Therefore, we implement $\mathbf{q}^{k}_{\epsilon}$ and estimate $\mathbf{\tilde{v}}(\mathbf{\dot{q}}^{k+1})$ based on it.
-This could still result in a minor temporary capacity limit violation, which should be manageable by the DER.
+This could still result in minor temporary DER capacity limit violations, which should be manageable by the DERs.
 To ensure strict capacity limit satisfaction, the original feasible set can be slightly deflated [14].
-The initial feasible $\mathbf{u}^{\tau}$ in the second step can be simply chosen as $\mathbf{q}^{k}$.
+The initial feasible $\mathbf{u}^{\tau}$ in the second step can be chosen as $\mathbf{q}^{k}$.
+Figure 2 shows the implemented reactive power setpoints and communication requirements in an outer iteration.
+Figure 2: Implemented reactive power setpoints and communication requirements in an outer iteration for the proposed strategy.
+II-E Convergence analysis
+The convergence analysis of the nested algorithm is based on the following two assumptions.
+First, we assume a linear relation holds between voltages and reactive power injections.
+That is, $\mathbf{v}=\mathbf{v}^{0}+\mathbf{X}\mathbf{q}$, where the constant term $\mathbf{v}^{0}$ captures the impact of other disturbances.
+Based on the first assumption, $\mathbf{v}(\mathbf{u}^{\tau})-\mathbf{v}(\mathbf{\dot{q}}^{k+1})$ is the exact gradient of the objective function in (4) at $\mathbf{u}^{\tau}$.
+The inner loop then represents a standard Euclidean gradient projection algorithm to solve the quadratic program (4).
+Consequently, $\mathbf{u}^{\tau}$} is a converging sequence by invoking Proposition 2.3.2 in [13] and $\alpha^{u}$ chosen with the rule therein.
+Second, we assume the inner iterations indeed converge after $T$ iterations.
+The overall nested algorithm thus represents a scaled PDGP algorithm to solve (1).
+By a transformation of variables defined by: $\mathbf{q}^{\prime}=\mathbf{X}^{\frac{1}{2}}\mathbf{q}$ and $\mathbf{v}^{\prime}=\gamma\mathbf{v}$ where $\gamma=\frac{\alpha^{d}}{\alpha}$ is the ratio between the dual and primal step sizes, the problem (1) can be cast as a quadratic program in the space of $\mathbf{q}^{\prime}$ and $\mathbf{v}^{\prime}$.
+Likewise, the overall nested algorithm can be cast as an equivalent unscaled PDGP algorithm in $\mathbf{q}^{\prime}$ and $\mathbf{v}^{\prime}$ [13].
+Invoking Theorem 3 in [5] and with $\alpha$ chosen sufficiently small within the upper bound defined therein, the overall algorithm then converges to the optimizer of the regularized saddle-point problem of (1) at a Q-linear rate.
+Remark: First, while the linear voltage-reactive power relation is assumed to synthesize the controller and analyze its convergence, AC power flow is run in the simulation to calculate the real voltages.
+Such an assumption was also made in the proofs of [1, 11].
+Second, a large $T$ is not necessary in the simulation.
+We use $T=4$ so the inner loop might not converge.
+However, the algorithm performs well in the simulation.
+Notably, a similar assumption was made in [1].
 III Simulation results
-In our simulation experiment, we compare the centralized OFO based on the gather-and-broadcast communication, the distributed two-metric approach in [8], and our proposed nested distributed approach in both static and dynamic cases.
-We study a 100% penetration scenario of photovoltaics (PVs) in a 96-node low-voltage distribution system, which experiences overvoltage issues.
-Details of the distribution system and the generation and load data with a 6-second resolution can be found in [11] and references therein.
+In the numerical simulation, we study a 100% penetration scenario of photovoltaics (PVs) in a 96-node system, which experiences overvoltage issues.
+Details of the system and the generation and load data with a 6-second resolution can be found in [7] and references therein.
 The PV inverters are assumed to be oversized by 20% to provide sufficient reactive power.
@@
 The unit for power is $kW/kVar$.
-The step sizes are chosen with a trial-and-error strategy [8] as $\alpha^{\lambda}=\alpha^{\mu}=10^{6}$, $\alpha=10^{-1}$ (for centralized), $\alpha=10^{-3}$ (for distributed), and $\alpha^{u}=10^{2}$.
-The exploration parameter $\epsilon$ is chosen as $10^{-5}$, resulting in a maximum DER reactive power limit violation of 2%.
-For each generation and load data point, an OFO iteration is run which yields tentative reactive power setpoints, followed by one exploration step and 5 inner feedback iterations of (5) ($T=5$).
-Figure 1: Voltage profiles over 200 iterations for two remote buses in the static case using the proposed and the two-metric approaches.
+The step sizes are chosen with a trial-and-error strategy [12] as $\alpha^{d}=10^{6}$, $\alpha=2\times 10^{-4}$, and $\alpha^{u}=10^{2}$.
+The exploration parameter $\epsilon$ is chosen as $10^{-5}$, resulting in a maximum DER reactive power limit violation of 1%.
+The primal and dual regularization is used to establish the theoretical convergence result.
+The impact of some small regularization parameters on the numerical results is negligible.
+In the simulation, we set them as $10^{-4}$.
+For each generation and load data point, an OFO iteration is run which yields tentative reactive power setpoints, followed by one exploration step and 4 inner feedback iterations of (5), i.e.
+we assume that every setpoint in Fig.
+2 is implemented for 1 second.
+Finally, the nested algorithm requires only basic arithmetic operations and is computationally very efficient.
+For a series implementation of a 4-hour simulation, i.e.
+in total 14400 iterations, it takes 7.6 seconds, averaging 0.5 millisecond per iteration.
+Figure 3: Voltage profiles over 1000 iterations for several buses in the static case using the centralized and the proposed approaches.
 We count both the outer and inner iterations for our proposed approach.
-Figure 2: Reactive power profiles after 200 iterations for the centralized, the proposed, and the two-metric approaches.
+Figure 4: Reactive power profiles after 1000 iterations for the centralized and the proposed approaches.
 Figure 5: Voltage profiles for the most remote bus in the dynamic simulation.
-The average voltage violation (AVV) values are $1.9\times 10^{-4}$, $1.3\times 10^{-4}$, and $4.6\times 10^{-3}$ pu for the centralized, the proposed, and the two-metric approaches, respectively, which is defined as $\operatorname{AVV}=\frac{1}{K}\sum_{k=1}^{K}[v^{k}-\overline{v}]^{+}$, where $k$ is the iteration index and $K$ is the number of iterations.
-In the static case, the generation and load data are kept unchanged to study the convergence of various approaches.
-Figure 1 shows that while the two-metric approach does not converge and fails to regulate voltages, our proposed approach converges quickly after approximately 154 iterations, which correspond to around 22 outer and 132 inner iterations.
-We also compare the final reached reactive power setpoints after 200 iterations.
-Figure 2 shows that the centralized and our proposed approaches converge to the same solution.
-This is reasonable since the gradient scaling only alters the trajectory, but does not change the solution to Problem (1).
-In the dynamic case, we study the performance of the approaches in a time-varying environment, where the approaches are implemented online without reaching convergence.
-Figure 3 shows the significant voltage limit violation without control and demonstrates that our proposed nested distributed approach successfully enforces the voltage limit even for the most remote, and therefore, most sensitive bus, achieving an average voltage violation (AVV) of $1.3\times 10^{-4}$ pu, which is better even than that of the centralized approach ($1.9\times 10^{-4}$ pu).
-We conjecture that this is because of the smoothing brought by the inner feedback loop.
-The two-metric approach, however, fails to regulate voltages, resulting in 34 times higher AVV.
+We define the average voltage violation (AVV) as $\operatorname{AVV}=\frac{1}{K}\sum_{k=1}^{K}[v^{k}-\overline{v}]^{+}$, where $k$ is the iteration index and $K$ is the number of iterations.
+The AVV values for the various approaches are: $1.5\times 10^{-2}$ (no control), $8.3\times 10^{-5}$ (centralized), $2.5\times 10^{-3}$ (two-metric), $2.1\times 10^{-4}$ (proposed), $3.3\times 10^{-3}$ (Tang 2024 [11]), $8.6\times 10^{-4}$ (Qu 2020 [12]).
+In the static case, the generation and load data are kept unchanged to study the convergence of the nested approach.
+Figure 3 shows that the proposed approach suppresses the initial voltage limit violations quickly and converges after around 300 total iterations, i.e.
+50 outer OFO iterations.
+The approach also utilizes the available grid capacity as efficiently as the centralized one.
+Figure 4 compares the converged reactive power setpoints, which shows that the centralized and the proposed approaches converge to the same solution.
+This is reasonable since the gradient scaling only alters the converging trajectory, but does not change the solution to (1).
+In the dynamic case, we study the performance of the approaches in a time-varying environment, where the approaches are implemented online to track the time-varying optimizer of (1) without waiting for them to converge.
+Figure 5 shows significant voltage limit violations without control and demonstrates that our proposed distributed nested feedback optimization approach successfully enforces the voltage limit even for the most remote, and therefore, the most sensitive bus.
+An average voltage violation (AVV), defined in Fig.
+5, of $2.1\times 10^{-4}$ pu is achieved, which is only higher than that of the centralized approach.
+The latter however requires a centralized communication architecture and is thus susceptible to single-point failures.
+The proposed nested feedback approach significantly outperforms other distributed ones, which is visible in Fig.
+5 and quantified using the AVV.
 IV Conclusion
-In this letter, we focused on the distributed voltage regulation problem in distribution systems.
-We showed that the previously proposed two-metric approach might fail.
-As a remedy, we proposed a nested feedback optimization approach.
-Our simulation results showed that the approach achieved even better voltage regulation than a centralized approach while only requiring short-range communication between physical neighbours.
-The proposed distributed approach however does not extend to congestion management while the centralized does so.
-The approach works for either active or reactive power control, but not joint control.
+In this paper, we focused on the distributed voltage regulation problem in distribution systems and proposed a nested feedback optimization approach.
+We theoretically analyzed its convergence.
+Our simulation results showed that the approach, while only requiring short-range communication between physical neighbours, achieved satisfactory voltage regulation and outperformed other distributed approaches.
+The proposed distributed approach does not extend to congestion management while the centralized does so.
+The current presentation is based on radial systems but the approach can be readily extended to meshed systems using the approach in [1] to construct the sensitivity matrix $\mathbf{X}$.
+Extending the approach to unbalanced systems is an interesting future direction, where the focus is to construct a positive definite $\mathbf{X}$ with a sparse inverse matrix.
+Finally, the approach works for either active or reactive power control, but not joint control.
 This is because the observation in Section II-D does not hold anymore since the voltage difference is no longer the gradient of either active or reactive power injection, but represents their joint non-separable impact.

### v3

@@
 Abstract
-We investigate the real-time distribution system voltage regulation problem employing online feedback optimization (OFO) and short-range communication between physical neighbours.
-OFO does not need an accurate grid topology nor estimated consumption of non-controllable loads, affords fast calculations, and demonstrates robustness to uncertainties due to its feedback-based nature.
-These characteristics render it particularly suitable for real-time distribution system applications.
-However, many OFO controllers require centralized communication, which lack robustness to single-point failures.
+We investigate the real-time voltage regulation problem in distribution systems employing online feedback optimization (OFO) with short-range communication between physical neighbours.
+OFO does not need an accurate grid model nor estimated consumption of non-controllable loads, affords fast calculations, and demonstrates robustness to uncertainties and disturbances, which render it particularly suitable for real-time distribution system applications.
+However, many OFO controllers require centralized communication, making them susceptible to single-point failures.
 This paper proposes a distributed OFO design based on a nested feedback optimization strategy and analyzes its convergence.
-Numerical results reveal that the proposed design achieves satisfactory voltage regulation and outperforms other distributed approaches.
+Numerical study results demonstrate that the proposed design achieves effective voltage regulation and outperforms other distributed and local approaches.
 Index Terms:
-Distributed communication, online feedback optimization, voltage regulation, reactive power control
+Distributed communication, online feedback optimization, reactive power control, voltage regulation
 I Introduction
-With the proliferation of distributed energy resources (DERs) in distribution systems, real-time voltage regulation has become a challenging task for distribution system operators (DSOs).
-While local control cannot guarantee the desired voltage regulation [1], model-based feedforward optimization requires an accurate grid model, load measurements, and lacks robustness to system uncertainties and disturbances [2].
-To overcome these limitations, online feedback optimization (OFO) has been proposed as a promising solution approach.
-OFO utilizes measurements as feedback and employs optimization algorithms as feedback controllers to steer physical systems towards their optimal operating points [2]; see Fig.
-1 for comparison between feedforward and feedback systems.
-The feedback-based implementation eliminates the need for an accurate grid model and real-time load data and brings robustness to OFO.
-Compared to deep reinforcement learning, OFO does not need a complicated training process and has guarantees to pursue grid-level objectives.
-Compared to using traditional voltage regulators such as on-load tap changers and capacitor banks, OFO provides fast voltage regulation and can complement these slow-reacting devices in real time [3].
-While many OFO controllers rely on centralized communication [4, 5, 6, 7, 8], distributed OFO based on communication between physical neighbours appears a compelling approach due to its robustness to single-point failures[9, 1, 3, 10, 11, 12].
-Among the existing distributed designs, [9, 1, 3] demand particular forms of objective functions, while [10] relies on disseminating global information through neighboring nodes.
+The rapid growth of distributed energy resources (DERs) has introduced significant uncertainty and volatility to distribution systems.
+With the high resistance/reactance ratios of distribution cables, voltage fluctuations are occurring more frequently.
+Consequently, real-time voltage regulation has become a significant operational challenge for distribution system operators (DSOs).
+The conventional approach of DSOs is grid reinforcement, but this demands significant investments, a skilled workforce, and considerable time.
+During this lengthy process, a complementary solution is to exploit the flexibility of network assets and DERs.
+In this context, online feedback optimization (OFO) has recently emerged as a promising strategy, e.g.
+in [1, 2, 3, 4, 5, 6].
+OFO utilizes measurements as feedback and employs optimization algorithms as feedback controllers to steer physical systems towards their optimal operating points [7], which are defined by the well-established optimal power flow (OPF) problem.
+Compared to directly solving the OPF problem in a feedforward manner, OFO does not need an accurate grid model nor data of non-controllable loads, affords fast calculations, and is robust against uncertainties and disturbances in distribution systems due to its feedback-based nature [8, 7, 9].
+Figure 1 illustrates feedforward and feedback systems.
+Figure 1: Block diagrams of feedforward and feedback systems for voltage regulation using reactive power, where $q$ represents system input (reactive power), $v$ represents system output (voltages), $\omega$ ($\widehat{\omega}$) represents (estimated) system disturbance (non-controllable loads), and $h$ maps system input and disturbance to system output.
+This figure is adapted from [10].
+OFO also brings distinct advantages over other existing approaches.
+Compared to local droop control [11, 3, 12], OFO ensures more reliable voltage regulation [3] and can pursue grid-level objectives.
+In contrast to deep reinforcement learning [13, 14], OFO does not need a complicated (often centralized offline) training process while offering theoretical guarantees to obtain globally optimal solutions.
+Furthermore, unlike using traditional voltage regulators such as on-load tap changers and capacitor banks [15], OFO provides rapid voltage regulation with inverter-interfaced DERs and can complement these slow-reacting devices in real time [16].
+While many OFO controllers rely on centralized communication [1, 17, 2, 8, 4, 5, 18, 9, 19, 20, 21], distributed OFO based on communication between physical neighbours appears to be a compelling approach due to its robustness to single-point failures[6, 3, 16, 22, 23, 24].
+Among the existing distributed designs, [6, 3, 16] demand particular forms of objective functions, while [22] relies on disseminating global information through neighbouring nodes.
 However, the systematic delays inherent in this information propagation may raise stability issues, potentially undermining its efficacy in practical distribution systems.
 In [23], the distributed design was achieved by only keeping components in the network sensitivity matrix related to local and neighbouring nodes.
-This however does not assure voltage regulation.
-In this paper, we propose a distributed OFO design related to that in [12], where the distributed implementation was enabled by the sparsity of the inverse of the network sensitivity matrix $\mathbf{X}$, i.e.
+This heuristic approach does not assure adequate voltage regulation.
+In this paper, we propose a distributed OFO design inspired by [24], where the distributed implementation was enabled by the sparsity of the inverse of the network sensitivity matrix $\mathbf{X}$, i.e.
 $\mathbf{X}^{-1}$ is sparse.
-Both approaches continue with scaling the gradient by $\mathbf{X}^{-1}$ followed by a projection step to ensure satisfaction of the DER capacity constraint.
+Both our proposed and their approaches continue with scaling the gradient by $\mathbf{X}^{-1}$ followed by a projection step to ensure satisfaction of the DER capacity constraint.
 It is conceptually attractive to use the Euclidean distance in the projection, which can be implemented using only local information.
-However, as shown in [13], the distance measured by the norm $\lVert\mathbf{z}\rVert_{\mathbf{X}}=\sqrt{\mathbf{z}^{\intercal}\mathbf{X}\mathbf{z}}$ should be used instead in the projection step to ensure descent iterations and thus algorithm convergence.
+However, as shown in [25], the distance measured by the norm $\lVert\mathbf{z}\rVert_{\mathbf{X}}=\sqrt{\mathbf{z}^{\intercal}\mathbf{X}\mathbf{z}}$ should be used instead in the projection step to ensure descent iterations and thus convergence.
 With the Euclidean distance, it would be a two-metric approach which can lead to algorithm divergence [25].
 Solving this non-Euclidean projection problem would however require global information since $\mathbf{X}$ is dense, contradicting the distributed OFO design.
-As a remedy, [12] introduced additional dualization of the DER capacity limits on top of the Euclidean projection to ensure the DER constraint satisfaction.
-Shown in their numerical result and our simulation, this strategy takes a prohibitively large number of iterations to converge, which jeopardizes its performance in online implementation.
-Figure 1: Block diagrams of feedforward and feedback systems for voltage regulation using reactive power, where $q$ represents system input (reactive power), $v$ represents system output (voltages), $\omega$ ($\widehat{\omega}$) represents (estimated) system disturbance (non-controllable loads), and $h$ maps system input and disturbance to system output.
+As a remedy, [24] introduced additional dualization of the DER capacity constraint on top of the Euclidean projection to ensure the DER constraint satisfaction.
+Nevertheless, this strategy takes a prohibitively large number of iterations to converge, which may jeopardize its performance in online implementation.
 We tackle this challenge by proposing an iterative algorithm to solve this non-Euclidean projection problem that requires only local information and operates in an online feedback-based manner.
-This leads to our overall nested online feedback optimization approach, where the outer loop comprises the OFO iterations requiring communication between physical neighbours and yielding tentative but not necessarily feasible DER setpoints, and the inner loop solves the non-Euclidean projection problem to map the tentative setpoints to actual feasible setpoints using local information.
-The remainder of this paper is structured as follows: Section II presents the problem formulation and the proposed nested approach and analyzes its convergence.
-Section III presents simulation results, while Section IV draws conclusions and discusses future work.
-II Online feedback optimization
-II-A Problem formulation
-Consider the distribution system voltage regulation problem using reactive power as formulated in (1).
-DERs are located in the set $\mathcal{N}$, and $q_{i}$ represents their respective reactive power injection.
+This leads to our overall nested feedback optimization approach, where the outer loop comprises the OFO iterations requiring communication between physical neighbours and yielding tentative but not necessarily feasible DER setpoints, while the inner loop solves the non-Euclidean projection problem to map the tentative setpoints to actual feasible setpoints using local information.
+To summarize, the main contributions of this paper are:
+  • A local iterative algorithm is proposed to solve the non-Euclidean projection problem that operates in an online feedback-based manner.
+  • An efficient nested feedback optimization strategy that ensures adequate voltage regulation while eliminating the need for an accurate grid model and consumption of non-controllable loads is developed.
+  • The proposed strategy is theoretically analyzed and numerically assessed on a 96-bus system with high-resolution generation and load data.
+The remainder of this paper is structured as follows: Section II presents the system modeling, problem formulation, and a centralized OFO strategy.
+Section III introduces the proposed nested approach and analyzes its convergence.
+Section IV presents a numerical study, while Section V draws conclusions and discusses future work.
+II Problem formulation and centralized OFO
+II-A Distribution system modeling
+[…]
+The linearized DistFlow equations in (1) were proposed in [26] and can be used to model balanced radial distribution systems.
+Equations (1a)-(1b) represent nodal active and reactive power balance constraints, respectively.
+The voltage relation is modeled in (1c), which further leverages the assumption that $v_{i}^{2}-v_{j}^{2}\approx 2(v_{i}-v_{j})$ since $v_{i}\approx 1\text{ pu},\forall i\in\mathcal{N}$.
+$P_{ij}+p_{j}=\sum_{k:(j,k)\in\mathcal{E}}P_{jk}+p^{d}_{j},\forall j\in\mathcal{N}^{+},$ (1a)
+$Q_{ij}+q_{j}=\sum_{k:(j,k)\in\mathcal{E}}Q_{jk}+q^{d}_{j},\forall j\in\mathcal{N}^{+},$ (1b)
+$v_{i}-v_{j}=r_{ij}P_{ij}+x_{ij}Q_{ij},\forall(i,j)\in\mathcal{E}.$ (1c)
+Applying (1) directly in a feedforward voltage regulation scheme requires estimates of $p_{j}^{d}$ and $q^{d}_{j}$, $\forall j\in\mathcal{N}^{+}$.
+Moreover, the modeling inaccuracy will also result in voltage approximation errors.
+To synthesize the proposed feedback controller for voltage regulation, we leverage (1) to derive network sensitivities, which relate changes in nodal active and reactive power injections to voltage variations.
+Following the steps in [27] or [28], the linear relation in (2) can be constructed.
+$\mathbf{v}=v_{0}\mathbf{1}+\mathbf{R}(\mathbf{p}-\mathbf{p}^{d})+\mathbf{X}(\mathbf{q}-\mathbf{q}^{d}).$ (2)
+The $N\times N$-dimensional symmetrical matrices $\mathbf{R}$ and $\mathbf{X}$ are given in (3), where $\mathcal{E}_{i}$ represents the set of cables on the unique path from the substation bus to bus $i$.
+$\mathbf{R}_{ij}=\sum_{(h,k)\in\mathcal{E}_{i}\cap\mathcal{E}_{j}}r_{hk},\mathbf{X}_{ij}=\sum_{(h,k)\in\mathcal{E}_{i}\cap\mathcal{E}_{j}}x_{hk}.$ (3)
+These matrices accordingly capture the network sensitivities, i.e.
+$\frac{\partial\mathbf{v}}{\partial\mathbf{p}}=\mathbf{R},\frac{\partial\mathbf{v}}{\partial\mathbf{q}}=\mathbf{X}.$ (4)
+When the resistances and reactances of the cables are all positive, [27] shows that $\mathbf{R}$ and $\mathbf{X}$ are both symmetrical positive definite.
+Furthermore, [24] shows that their inverse matrices $\mathbf{R}^{-1}$ and $\mathbf{X}^{-1}$ have the sparsity pattern such that
+$\mathbf{R}_{ij}^{-1}\neq 0,\mathbf{X}_{ij}^{-1}\neq 0\iff(i,j)\in\mathcal{E}\text{ or }i=j.$ (5)
+Figure 2 visualizes $\mathbf{X}$ and $\mathbf{X}^{-1}$ for the test system in Section IV, where the sparsity patterns are shown.
+(a) $\mathbf{X}$ (pu/kVar)
+(b) $\mathbf{X}^{-1}$ (kVar/pu)
+Figure 2: Heat maps of the network sensitivity matrix $\mathbf{X}$ and its inverse $\mathbf{X}^{-1}$.
+II-B Problem formulation
+Consider the problem of distribution system voltage regulation using reactive power as formulated in (6).
 The objective function (6a) minimizes the reactive power use cost, where $c_{i}>0$ is the cost parameter.
 Constraint (6b) enforces the lower and upper voltage limits $\underline{v}$ and $\overline{v}$, where $\mu_{i}$ and $\lambda_{i}$ are the respective dual variables.
-Finally, (1c) ensures that the reactive power injection $q_{i}$ respects the lower and upper limits $\underline{q}_{i}$ and $\overline{q}_{i}$.
-Note that $v_{i}$ is dependent on $q_{i}$ and other disturbances in the system such as active and reactive power consumption of non-controllable load units.
-$\underset{q_{i},\forall i\in\mathcal{N}}{\operatorname{minimize}}$ $\quad f:=\frac{1}{2}c_{i}q_{i}^{2}$ (1a)
-$\operatorname{s.t.}$ $\quad\underline{v}\leq v_{i}\leq\overline{v}:\mu_{i},\lambda_{i},\forall i\in\mathcal{N}$ (1b)
-$\underline{q}_{i}\leq q_{i}\leq\overline{q}_{i},\forall i\in\mathcal{N}$ (1c)
-II-B Centralized primal-dual gradient projection algorithm
-Solving (1) offline in a feedforward way demands an accurate grid model and disturbance data to evaluate $v_{i}$, which might not be available.
+Finally, (6c) ensures that the reactive power injection $q_{i}$ respects the lower and upper limits $\underline{q}_{i}$ and $\overline{q}_{i}$, which capture DER capacity and power factor constraints.
+It is noteworthy that $v_{i}$ is dependent on the decision variable $q_{i}$ and other disturbances in the system such as active and reactive power consumption of non-controllable loads.
+$\underset{q_{i},\forall i\in\mathcal{N}^{+}}{\operatorname{minimize}}$ $\quad\sum_{i\in\mathcal{N}^{+}}\frac{1}{2}c_{i}q_{i}^{2},$ (6a)
+$\operatorname{s.t.}$ $\quad\underline{v}\leq v_{i}\leq\overline{v}:\mu_{i},\lambda_{i},\forall i\in\mathcal{N}^{+},$ (6b)
+$\underline{q}_{i}\leq q_{i}\leq\overline{q}_{i},\forall i\in\mathcal{N}^{+}.$ (6c)
+For further developments, the dual problem of (6) is formulated in (7).
+$\underset{\mu_{i}\geq 0,\lambda_{i}\geq 0,\forall i\in\mathcal{N}^{+}}{\operatorname{maximize}}\big{\{}\underset{\underline{q}_{i}\leq q_{i}\leq\overline{q}_{i},,\forall i\in\mathcal{N}^{+}}{\operatorname{minimize}}\quad\mathcal{L}(\mathbf{q},\boldsymbol{\mu},\boldsymbol{\lambda})\big{\}},$ (7)
+where the partial Lagrangian function is defined in (8).
+$\mathcal{L}(\mathbf{q},\boldsymbol{\mu},\boldsymbol{\lambda})=\sum_{i\in\mathcal{N}^{+}}\left[\frac{1}{2}c_{i}q_{i}^{2}+\lambda_{i}\left(v_{i}-\overline{v}\right)+\mu_{i}\left(\underline{v}-v_{i}\right)\right].$ (8)
+II-C Centralized primal-dual gradient projection algorithm
+Solving (6) offline in a feedforward way demands an accurate grid model and disturbance data to evaluate $v_{i}$, which might not be readily available.
 The feedforward approach also lacks robustness [7].
-Thus, this section introduces an instance of OFO based on the primal-dual gradient projection (PDGP) algorithm [4, 5] to track the time-varying optimizer of (1), which requires a centralized gather-and-broadcast communication architecture.
-The PDGP algorithm performs projected gradient ascent and descent steps iteratively for the dual and primal variables, respectively.
+Thus, this section introduces a feedback controller based on OFO.
+The controller leverages primal-dual gradient projection (PDGP) as the underlying algorithm [1, 17, 2] to track the time-varying optimizers of (6), which requires a centralized gather-and-broadcast communication architecture.
+Specifically, the PDGP algorithm performs projected gradient ascent and descent steps iteratively for the dual and primal variables, respectively.
 At each time step (iteration) $k$, it repeats the following three steps:
   1. 1.
-    For each bus $i\in\mathcal{N}$, collect its voltage measurement $\tilde{v}_{i}^{k}$ and update $\lambda_{i}$ and $\mu_{i}$, where the projection operator is defined as $[u]^{+}:=\max(u,0)$, $\alpha$ with various superscripts represents step sizes, while $r^{d}$ and $r^{p}$ in (2c) are the dual and primal regularization factors [4, 5], respectively.
-    $\lambda_{i}^{k+1}\leftarrow\big{[}\lambda_{i}^{k}+\alpha^{d}(\tilde{v}_{i}^{k}-\overline{v}-r^{d}\lambda_{i}^{k})\big{]}^{+}$ (2a)
-    $\mu_{i}^{k+1}\leftarrow\big{[}\mu_{i}^{k}+\alpha^{d}(\underline{v}-\tilde{v}_{i}^{k}-r^{d}\mu_{i}^{k})\big{]}^{+}$ (2b)
+    For each bus $i\in\mathcal{N}^{+}$, collect its voltage measurement $\tilde{v}_{i}^{k}$ and update $\lambda_{i}$ and $\mu_{i}$ using (9a)-(9b), where the projection operator is defined as $[u]^{+}=\max(u,0)$, $\alpha$ with various superscripts represents step sizes, while $r^{d}$ and $r^{p}$ in (9c) are the dual and primal regularization factors [1, 17, 2], respectively.
+    $\lambda_{i}^{k+1}=\big{[}\lambda_{i}^{k}+\alpha^{d}(\tilde{v}_{i}^{k}-\overline{v}-r^{d}\lambda_{i}^{k})\big{]}^{+},$ (9a)
+    $\mu_{i}^{k+1}=\big{[}\mu_{i}^{k}+\alpha^{d}(\underline{v}-\tilde{v}_{i}^{k}-r^{d}\mu_{i}^{k})\big{]}^{+}.$ (9b)
   2. 2.
-    The DSO broadcasts the updated dual variables $\boldsymbol{\lambda}^{k+1}$ and $\boldsymbol{\mu}^{k+1}$ to all buses, where $\boldsymbol{\lambda}^{k+1}$ and $\boldsymbol{\mu}^{k+1}$ are vectors constructed by stacking $\lambda_{i}^{k+1}$ and $\mu_{i}^{k+1}$, respectively.
+    The DSO broadcasts the updated dual variables $\boldsymbol{\lambda}^{k+1}$ and $\boldsymbol{\mu}^{k+1}$ to all buses.
   3. 3.
-    For each bus $i\in\mathcal{N}$, update its reactive power setpoints $q_{i}$, where $[u]_{\underline{q}_{i}}^{\overline{q}_{i}}$ represents the Euclidean projection into $[\underline{q}_{i},\overline{q}_{i}]$, $\mathbf{X}$ is a $|\mathcal{N}|$-dimensional dense matrix mapping reactive power injections to voltages, i.e.
-the sensitivity matrix, as constructed in [12], and $\mathbf{q}^{k}$ is the stack of $q^{k}_{i}$.
+    For each bus $i\in\mathcal{N}^{+}$, update its reactive power setpoints $q_{i}$ using (9c), where $[u]_{\underline{u}}^{\overline{u}}$ represents the Euclidean projection into $[\underline{u},\overline{u}]$.
 Since $\mathbf{X}$ is dense, Step 2 is needed to provide information for Step 3.
-    $q_{i}^{k+1}\leftarrow\big{[}q_{i}^{k}-\alpha[c_{i}q_{i}^{k}+[\mathbf{X}(\boldsymbol{\lambda}^{k+1}-\boldsymbol{\mu}^{k+1}+r^{p}\mathbf{q}^{k})]_{i}]\big{]}_{\underline{q}_{i}^{k}}^{\overline{q}_{i}^{k}}$ (2c)
-II-C Two-metric distributed approach
-To avoid the centralized communication requirement and enable the distributed communication architecture where communication is only between physical neighbours, [12] proposed to scale the gradient in (2c) by a sparse positive definite matrix $\mathbf{X}^{-1}$, which has non-zero entries only in the diagonal and the $i$-th row $j$-th column if buses $i$ and $j$ are adjacent.
-Continuing with the Euclidean projection yields (3) which replaces (2c) in the OFO where $\mathbf{C}$ is a $|\mathcal{N}|$-dimensional diagonal matrix with $c_{i}$ defined in (1a).
-$q_{i}^{k+1}\leftarrow\big{[}q_{i}^{k}-\alpha[[\mathbf{X}^{-1}\mathbf{C}\mathbf{q}^{k}]_{i}+\lambda_{i}^{k+1}-\mu_{i}^{k+1}+r^{p}q_{i}^{k}]\big{]}_{\underline{q}_{i}^{k}}^{\overline{q}_{i}^{k}}$ (3)
+    $q_{i}^{k+1}=\big{[}q_{i}^{k}-\alpha[c_{i}q_{i}^{k}+[\mathbf{X}(\boldsymbol{\lambda}^{k+1}-\boldsymbol{\mu}^{k+1}+r^{p}\mathbf{q}^{k})]_{i}]\big{]}_{\underline{q}_{i}^{k}}^{\overline{q}_{i}^{k}}.$ (9c)
+III Distributed OFO
+III-A Two-metric distributed approach
+To avoid the centralized communication requirement and enable the distributed communication architecture where communication is only between physical neighbours, [24] proposed to scale the gradient in (9c) by the sparse positive definite matrix $\mathbf{X}^{-1}$, which has non-zero entries only in the diagonal and the $i$-th row $j$-th column if buses $i$ and $j$ are adjacent, i.e.
+as in (5).
+Continuing with the Euclidean projection yields (10) which replaces (9c) in the OFO, where $\mathbf{C}$ is a $N$-dimensional diagonal matrix with $c_{i}$ defined in (6a).
+$q_{i}^{k+1}=\big{[}q_{i}^{k}-\alpha[[\mathbf{X}^{-1}\mathbf{C}\mathbf{q}^{k}]_{i}+\lambda_{i}^{k+1}-\mu_{i}^{k+1}+r^{p}q_{i}^{k}]\big{]}_{\underline{q}_{i}^{k}}^{\overline{q}_{i}^{k}}.$ (10)
 Computing (10) for node $i$ now requires only information exchange with its physical neighbours.
-We stack $q_{i}^{k}-\alpha[[\mathbf{X}^{-1}\mathbf{C}\mathbf{q}^{k}]_{i}+\lambda_{i}^{k+1}-\mu_{i}^{k+1}+r^{p}q_{i}^{k}]$ for all nodes (tentative but not necessarily feasible setpoints to be projected) as the vector $\mathbf{\dot{q}}^{k+1}$ and define $\mathcal{X}^{k}$ as $[\underline{q}_{1}^{k},\overline{q}_{1}^{k}]\times[\underline{q}_{2}^{k},\overline{q}_{2}^{k}]\times\cdots\times[\underline{q}_{|\mathcal{N}|}^{k},\overline{q}_{|\mathcal{N}|}^{k}]$, where $\times$ is the Cartesian product of sets.
-II-D Proposed distributed approach
-The two-metric approach in Section II-C does not necessarily lead to algorithm convergence since it is not in general a descent iteration [13].
-To ensure convergence, the projection in (3) should use the distance measured in terms of the norm $\lVert\mathbf{z}\rVert_{\mathbf{X}}=\sqrt{\mathbf{z}^{\intercal}\mathbf{X}\mathbf{z}}$ due to the scaling matrix $\mathbf{X}^{-1}$ [13], that is:
-$\mathbf{q}^{k+1}\leftarrow\underset{\mathbf{u}\in\mathcal{X}^{k}}{\operatorname{argmin}}\quad\frac{1}{2}(\mathbf{u}-\mathbf{\dot{q}}^{k+1})^{\intercal}\mathbf{X}(\mathbf{u}-\mathbf{\dot{q}}^{k+1})$ (4)
-This however contradicts the distributed communication architecture of the overall OFO algorithm since, unlike the Euclidean projection (equivalent to the case where $\mathbf{X}$ is the identity matrix), solving (4) again requires global information.
+However, the two-metric approach does not necessarily lead to algorithm convergence since it is not in general a descent iteration [25].
+This is illustrated in Fig.
+3 with an example.
+For further developments, we define
+$\mathbf{\dot{q}}^{k+1}=$ (11)
+$\quad\left[q_{i}^{k}-\alpha[[\mathbf{X}^{-1}\mathbf{C}\mathbf{q}^{k}]_{i}+\lambda_{i}^{k+1}-\mu_{i}^{k+1}+r^{p}q_{i}^{k}],i\in\mathcal{N}^{+}\right],$
+which are tentative but not necessarily feasible setpoints to be projected for all nodes, and define
+$\mathcal{X}^{k}=[\underline{q}_{1}^{k},\overline{q}_{1}^{k}]\times[\underline{q}_{2}^{k},\overline{q}_{2}^{k}]\times\cdots\times[\underline{q}_{N}^{k},\overline{q}_{N}^{k}],$ (12)
+where $\times$ is the Cartesian product of sets.
+Figure 3: An illustration showing that the two-metric approach is not a descent iteration, where $\mathbf{x}^{k}$ is the current iterate which is also the optimum, $X$ is the feasible region, and $\nabla f^{k}$ is the gradient at $\mathbf{x}^{k}$.
+With (plain) gradient projection on the left panel, the iterate stays at the optimum, i.e.
+$\mathbf{x}^{k+1}=\mathbf{x}^{k}$.
+With the two-metric approach on the right panel where $\mathbf{Y}$ represents the gradient scaling metric, $\mathbf{x}^{k+1}$ leaves the optimum.
+III-B Proposed distributed approach
+It has been shown in [25] that, to ensure convergence, the projection should use the distance measured in terms of the norm $\lVert\mathbf{z}\rVert_{\mathbf{X}}=\sqrt{\mathbf{z}^{\intercal}\mathbf{X}\mathbf{z}}$ due to the scaling matrix $\mathbf{X}^{-1}$ [25], that is:
+$\mathbf{q}^{k+1}=\underset{\mathbf{u}\in\mathcal{X}^{k}}{\operatorname{argmin}}\quad\frac{1}{2}(\mathbf{u}-\mathbf{\dot{q}}^{k+1})^{\intercal}\mathbf{X}(\mathbf{u}-\mathbf{\dot{q}}^{k+1}).$ (13)
+This however contradicts the distributed communication architecture of the overall OFO algorithm since, unlike the Euclidean projection, solving (13) again requires global information.
 As a solution, [24] proposed dualization of the DER capacity limits, combined with the Euclidean projection, to ensure the DER constraint satisfaction.
-However, as demonstrated both in their numerical result and our simulation, this approach requires an excessively large number of iterations to converge, significantly hindering its performance in online implementation.
+However, this approach requires an excessively large number of iterations to reach convergence, significantly hindering its performance in online implementation.
+Figure 4: Illustration of the approximation.
 To tackle this challenge, this paper proposes an iterative algorithm to solve (13) that requires only local information and is implemented in an online feedback-based manner.
 This leads to our nested feedback optimization approach, where the outer loop comprises the OFO iterations giving tentative reactive power setpoints, and the inner loop solves the non-Euclidean projection problem (13) to generate feasible setpoints.
-Our key observation is that given a feasible $\mathbf{u}^{\tau}\in\mathcal{X}^{k}$, the gradient of the objective function in (4) at $\mathbf{u}^{\tau}$ is $\mathbf{X}(\mathbf{u}^{\tau}-\mathbf{\dot{q}}^{k+1})$, which can be approximated by the voltage difference $\mathbf{v}(\mathbf{u}^{\tau})-\mathbf{v}(\mathbf{\dot{q}}^{k+1})$ assuming system disturbances do not change significantly when all nodes are locally implementing this inner iterative algorithm.
+Our key observation is that given a feasible $\mathbf{u}^{\tau}\in\mathcal{X}^{k}$, the gradient of the objective function in (13) at $\mathbf{u}^{\tau}$ is $\mathbf{X}(\mathbf{u}^{\tau}-\mathbf{\dot{q}}^{k+1})$, which can be approximated by the voltage difference $\mathbf{v}(\mathbf{u}^{\tau})-\mathbf{v}(\mathbf{\dot{q}}^{k+1})$ according to (2) assuming system disturbances do not change significantly when all nodes are locally implementing this inner iterative algorithm, i.e.
+$\mathbf{X}(\mathbf{u}^{\tau}-\mathbf{\dot{q}}^{k+1})\approx\mathbf{v}(\mathbf{u}^{\tau})-\mathbf{v}(\mathbf{\dot{q}}^{k+1}).$ (14)
 The proposed algorithm is based on the non-scaled Euclidean gradient projection and proceeds as:
   1. 1.
-    Implement $\mathbf{q}^{k}_{\epsilon}:=\mathbf{q}^{k}+\epsilon(\mathbf{\dot{q}}^{k+1}-\mathbf{q}^{k})$ and measure the voltage $\mathbf{\tilde{v}}(\mathbf{q}^{k}_{\epsilon})$.
-Estimate $\mathbf{\tilde{v}}(\mathbf{\dot{q}}^{k+1})$ as $\mathbf{\tilde{v}}^{k}$ + ($\mathbf{\tilde{v}}(\mathbf{q}^{k}_{\epsilon})$-$\mathbf{\tilde{v}}^{k}$)/$\epsilon$.
+    Implement
+    $\mathbf{q}^{k}_{\epsilon}=\mathbf{q}^{k}+\epsilon(\mathbf{\dot{q}}^{k+1}-\mathbf{q}^{k}),$ (15)
+    and measure the voltage $\mathbf{\tilde{v}}(\mathbf{q}^{k}_{\epsilon})$. Estimate
+    $\mathbf{\tilde{v}}(\mathbf{\dot{q}}^{k+1})\approx\mathbf{\tilde{v}}^{k}+(\mathbf{\tilde{v}}(\mathbf{q}^{k}_{\epsilon})-\mathbf{\tilde{v}}^{k})/\epsilon.$ (16)
   2. 2.
     Given a feasible $\mathbf{u}^{\tau}$, implement it, measure the voltage $\mathbf{\tilde{v}}(\mathbf{u}^{\tau})$, compute the next implementable reactive power setpoint $\mathbf{u}^{\tau+1}$ as (17), and repeat this step for $T$ iterations.
-    $\mathbf{u}^{\tau+1}\leftarrow[\mathbf{u}^{\tau}-\alpha^{u}(\mathbf{\tilde{v}}(\mathbf{u}^{\tau})-\mathbf{\tilde{v}}(\mathbf{\dot{q}}^{k+1}))]_{\mathbf{\underline{q}}^{k}}^{\mathbf{\overline{q}}^{k}}$ (5)
-In the first step, $\epsilon$ represents a small exploration parameter, which is similar to that in model-free feedback optimization [14].
+    $\mathbf{u}^{\tau+1}=[\mathbf{u}^{\tau}-\alpha^{u}(\mathbf{\tilde{v}}(\mathbf{u}^{\tau})-\mathbf{\tilde{v}}(\mathbf{\dot{q}}^{k+1}))]_{\mathbf{\underline{q}}^{k}}^{\mathbf{\overline{q}}^{k}}.$ (17)
+Remark 1.
+In the first step, $\epsilon$ represents a small exploration parameter, which is similar to that in model-free feedback optimization [29, 30, 31, 32].
 Since $\mathbf{\dot{q}}^{k+1}$ can be pretty far away from the feasible set, it might be impossible to implement it.
 Therefore, we implement $\mathbf{q}^{k}_{\epsilon}$ and estimate $\mathbf{\tilde{v}}(\mathbf{\dot{q}}^{k+1})$ based on it.
+This is illustrated in Fig.
+4.
 This could still result in minor temporary DER capacity limit violations, which should be manageable by the DERs.
@@
 The initial feasible $\mathbf{u}^{\tau}$ in the second step can be chosen as $\mathbf{q}^{k}$.
-Figure 2 shows the implemented reactive power setpoints and communication requirements in an outer iteration.
-Figure 2: Implemented reactive power setpoints and communication requirements in an outer iteration for the proposed strategy.
-II-E Convergence analysis
+The online implementation of the overall nested algorithm is summarized in Algorithm 1.
+Data: $\mathbf{q}^{0}$
+Result: $\mathbf{q}^{k}$, $k\geq 0$; $\mathbf{u}^{\tau}$, $\tau\leq T$
+1 Set k = 0;
+2 while $k\geq 0$ do
+3       Implement $\mathbf{q}^{k}$ and collect voltage measurement $\tilde{\mathbf{v}}(\mathbf{q}^{k})$;
+4       Communicate $\mathbf{q}^{k}$ between physical neighbours and calculate $\dot{\mathbf{q}}^{k+1}$ via (11);
+5       Implement $\mathbf{q}^{k}_{\epsilon}$, measure voltage $\mathbf{\tilde{v}}(\mathbf{q}^{k}_{\epsilon})$ and estimate $\mathbf{\tilde{v}}(\mathbf{\dot{q}}^{k+1})$;
+6       Set $\tau=0$;
+7       while $\tau<T$ do
+8             Implement $\mathbf{u}^{\tau}$ and measure voltage $\tilde{\mathbf{v}}(\mathbf{u}^{\tau})$ (choose $\mathbf{u}^{0}$ as $\mathbf{q}^{k}$);
+9             Calculate next implementable $\mathbf{u}^{\tau+1}$ via (17);
+10             Set $\tau=\tau+1$;
+11
+12       end while
+13      Set $\mathbf{q}^{k+1}=\mathbf{u}^{T}$;
+14       Set $k=k+1$.
+15 end while
+Algorithm 1 Nested feedback optimization algorithm
+III-C Convergence analysis
 The convergence analysis of the nested algorithm is based on the following two assumptions.
-First, we assume a linear relation holds between voltages and reactive power injections.
-That is, $\mathbf{v}=\mathbf{v}^{0}+\mathbf{X}\mathbf{q}$, where the constant term $\mathbf{v}^{0}$ captures the impact of other disturbances.
+Assumption 1.
+The linear relation (2) holds between voltages and reactive power injections.
 Based on the first assumption, $\mathbf{v}(\mathbf{u}^{\tau})-\mathbf{v}(\mathbf{\dot{q}}^{k+1})$ is the exact gradient of the objective function in (13) at $\mathbf{u}^{\tau}$.
 The inner loop then represents a standard Euclidean gradient projection algorithm to solve the quadratic program (13).
-Consequently, $\mathbf{u}^{\tau}$} is a converging sequence by invoking Proposition 2.3.2 in [13] and $\alpha^{u}$ chosen with the rule therein.
-Second, we assume the inner iterations indeed converge after $T$ iterations.
+Consequently, $\mathbf{u}^{\tau}$} is a converging sequence by invoking Proposition 2.3.2 in [25] with $0<\alpha^{u}<2/\lambda_{max}(\mathbf{X})$ where $\lambda_{max}(\mathbf{X})$ is the largest eigenvalue of $\mathbf{X}$.
+Assumption 2.
+The inner iterations reach convergence after $T$ iterations.
 The overall nested algorithm thus represents a scaled PDGP algorithm to solve (6).
-By a transformation of variables defined by: $\mathbf{q}^{\prime}=\mathbf{X}^{\frac{1}{2}}\mathbf{q}$ and $\mathbf{v}^{\prime}=\gamma\mathbf{v}$ where $\gamma=\frac{\alpha^{d}}{\alpha}$ is the ratio between the dual and primal step sizes, the problem (1) can be cast as a quadratic program in the space of $\mathbf{q}^{\prime}$ and $\mathbf{v}^{\prime}$.
-Likewise, the overall nested algorithm can be cast as an equivalent unscaled PDGP algorithm in $\mathbf{q}^{\prime}$ and $\mathbf{v}^{\prime}$ [13].
-Invoking Theorem 3 in [5] and with $\alpha$ chosen sufficiently small within the upper bound defined therein, the overall algorithm then converges to the optimizer of the regularized saddle-point problem of (1) at a Q-linear rate.
-Remark: First, while the linear voltage-reactive power relation is assumed to synthesize the controller and analyze its convergence, AC power flow is run in the simulation to calculate the real voltages.
+By a transformation of variables defined by: $\mathbf{q}^{\prime}=\mathbf{X}^{\frac{1}{2}}\mathbf{q}$ and $\mathbf{v}^{\prime}=\gamma\mathbf{v}$ where $\gamma=\sqrt{\frac{\alpha^{d}}{\alpha}}$ is the squared root of the ratio between the dual and primal step sizes, the problem (6) can be cast as an equivalent quadratic program in the space of $\mathbf{q}^{\prime}$ and $\mathbf{v}^{\prime}$ as in (18).
+$\underset{\mathbf{q^{\prime}}}{\operatorname{minimize}}$ $\quad\frac{1}{2}\mathbf{q^{\prime}}^{\intercal}\mathbf{X}^{-\frac{1}{2}}\mathbf{C}\mathbf{X}^{-\frac{1}{2}}\mathbf{q^{\prime}},$ (18a)
+$\operatorname{s.t.}$ $\quad\gamma\mathbf{\underline{v}}\leq\mathbf{v^{\prime}}\leq\gamma\mathbf{\overline{v}}:\boldsymbol{\mu^{\prime}},\boldsymbol{\lambda^{\prime}},$ (18b)
+$\mathbf{\underline{q}}\leq\mathbf{X}^{-\frac{1}{2}}\mathbf{q^{\prime}}\leq\mathbf{\overline{q}}.$ (18c)
+The unscaled regularized PDGP algorithm to solve (18) includes the following iterations:
+$\boldsymbol{\lambda^{\prime}}^{k+1}=[\boldsymbol{\lambda^{\prime}}^{k}+\alpha(\gamma\mathbf{\tilde{v}}-\gamma\mathbf{\overline{v}}-r^{d}\boldsymbol{\lambda^{\prime}}^{k})]^{+},$ (19a)
+$\boldsymbol{\mu^{\prime}}^{k+1}=[\boldsymbol{\mu^{\prime}}^{k}+\alpha(\gamma\mathbf{\underline{v}}-\gamma\mathbf{\tilde{v}}-r^{d}\boldsymbol{\mu^{\prime}}^{k})]^{+},$ (19b)
+$\mathbf{q^{\prime}}^{k+1}=\underset{\mathbf{q^{\prime},\mathbf{\underline{q}}\leq\mathbf{X}^{-\frac{1}{2}}\mathbf{q^{\prime}}\leq\mathbf{\overline{q}}}}{\operatorname{argmin}}\lVert\mathbf{q^{\prime}}-\big{[}\mathbf{q^{\prime}}^{k}-\alpha[\mathbf{X}^{-\frac{1}{2}}$ (19c)
+$\qquad\cdot\mathbf{C}\mathbf{X}^{-\frac{1}{2}}\mathbf{q^{\prime}}^{k}+\gamma\mathbf{X}^{\frac{1}{2}}(\boldsymbol{\lambda^{\prime}}^{k+1}-\boldsymbol{\mu^{\prime}}^{k+1})+r^{p}\mathbf{q^{\prime}}^{k}]\big{]}\rVert_{2}^{2}.$
+which are equivalent to (9a)-(9b) and (11)-(13) [25, Eqs.
+(2.37) and (2.38)].
+Invoking Theorem 3 in [2] with $\alpha$ chosen using [2, Eq.
+(34)], (19) then converges to the optimizer of the regularized saddle-point problem of (18).
+Consequently, our proposed nested algorithm also converges to the optimizer of the regularized saddle-point problem of (6).
+Remark 2.
+First, while the linear voltage-reactive power relation is assumed to synthesize the controller and analyze its convergence, AC power flow is run in the simulation to calculate the actual voltages.
 Such an assumption was also made in the proofs of [3, 23].
@@
 Notably, a similar assumption was made in [3].
-III Simulation results
-In the numerical simulation, we study a 100% penetration scenario of photovoltaics (PVs) in a 96-node system, which experiences overvoltage issues.
-Details of the system and the generation and load data with a 6-second resolution can be found in [7] and references therein.
+IV Case study
+IV-A Case description
+Figure 5: A 96-bus test system from Simbench [33].
+In the numerical simulation, we study a 100% penetration scenario of photovoltaics (PVs) in a 96-bus system, i.e.
+we assume every node has a PV installation.
+The test system is adopted from Simbench [33] and is shown in Fig.
+5, which is expected to experience overvoltage issues.
+The network topology and impedance data are the same as those in the original dataset.
+For each bus, a PV installation with a randomly generated DC capacity from 3-10 kW is assumed.
 The PV inverters are assumed to be oversized by 20% to provide sufficient reactive power.
+For the dynamic case study, generation and load data with a 6-second resolution are used.
+The PV profiles are linearly interpolated using the HelioClim-3 dataset [34] with a 1-minute resolution.
+Base load profiles are obtained by aggregating 1-second resolution data from the ECO dataset [35].
+AC power flow problems are solved with a high-performance library PowerGridModel [36].
 The reactive power cost matrix $\mathbf{C}$ is chosen as the identity matrix.
@@
 The unit for power is kW/kVar.
-The step sizes are chosen with a trial-and-error strategy [12] as $\alpha^{d}=10^{6}$, $\alpha=2\times 10^{-4}$, and $\alpha^{u}=10^{2}$.
-The exploration parameter $\epsilon$ is chosen as $10^{-5}$, resulting in a maximum DER reactive power limit violation of 1%.
+The step sizes are chosen with a trial-and-error strategy [24] as $\alpha^{d}=10^{6}$, $\alpha=5\times 10^{-4}$, and $\alpha^{u}=10^{2}$.
+The exploration parameter $\epsilon$ is chosen as $10^{-5}$, resulting in a maximum DER reactive power limit violation of 0.6%.
 The primal and dual regularization is used to establish the theoretical convergence result.
-The impact of some small regularization parameters on the numerical results is negligible.
+The impact of some small regularization parameters on numerical results is negligible.
 In the simulation, we set them as $10^{-4}$.
 For each generation and load data point, an OFO iteration is run which yields tentative reactive power setpoints, followed by one exploration step and 4 inner feedback iterations of (17), i.e.
-we assume that every setpoint in Fig.
-2 is implemented for 1 second.
+we assume that every setpoint in Algorithm 1 is implemented for 1 second.
 Finally, the nested algorithm requires only basic arithmetic operations and is computationally very efficient.
 For a series implementation of a 4-hour simulation, i.e.
-in total 14400 iterations, it takes 7.6 seconds, averaging 0.5 millisecond per iteration.
-Figure 3: Voltage profiles over 1000 iterations for several buses in the static case using the centralized and the proposed approaches.
+in total 14400 iterations, it takes 8.1 seconds, averaging 0.56 milliseconds per iteration.
+Figure 6: Voltage profiles over 500 iterations for several buses in the static case using the centralized and the proposed approaches.
 We count both the outer and inner iterations for our proposed approach.
-Figure 4: Reactive power profiles after 1000 iterations for the centralized and the proposed approaches.
-Figure 5: Voltage profiles for the most remote bus in the dynamic simulation.
-We define the average voltage violation (AVV) as $\operatorname{AVV}=\frac{1}{K}\sum_{k=1}^{K}[v^{k}-\overline{v}]^{+}$, where $k$ is the iteration index and $K$ is the number of iterations.
-The AVV values for the various approaches are: $1.5\times 10^{-2}$ (no control), $8.3\times 10^{-5}$ (centralized), $2.5\times 10^{-3}$ (two-metric), $2.1\times 10^{-4}$ (proposed), $3.3\times 10^{-3}$ (Tang 2024 [11]), $8.6\times 10^{-4}$ (Qu 2020 [12]).
+Figure 7: Reactive power profiles after 500 iterations for the centralized and the proposed approaches.
+(a) No control
+(b) Centralized
+(c) Proposed
+(d) Two metric
+(e) Tang 2024 [23]
+(f) Qu 2020 [24]
+(g) Droop control [37, 38]
+(h) Summary for bus 95
+Figure 8: Voltage profiles under different approaches in the dynamic simulation.
+(a) Centralized
+(b) Proposed
+(c) Two metric
+(d) Tang 2024 [23]
+(e) Qu 2020 [24]
+(f) Droop control [37, 38]
+Figure 9: Reactive power profiles under different approaches in the dynamic simulation.
+IV-B Static case results
 In the static case, the generation and load data are kept unchanged to study the convergence of the nested approach.
-Figure 3 shows that the proposed approach suppresses the initial voltage limit violations quickly and converges after around 300 total iterations, i.e.
+The simulation uses data at 12:00.
+Figure 6 shows that the proposed approach suppresses the initial voltage limit violations quickly and converges after around 200 total iterations, i.e.
 33 outer OFO iterations.
-The approach also utilizes the available grid capacity as efficiently as the centralized one.
+The approach also utilizes the available voltage limit as efficiently as the centralized one.
 Figure 7 compares the converged reactive power setpoints, which shows that the centralized and the proposed approaches converge to the same solution.
 This is reasonable since the gradient scaling only alters the converging trajectory, but does not change the solution to (6).
-In the dynamic case, we study the performance of the approaches in a time-varying environment, where the approaches are implemented online to track the time-varying optimizer of (1) without waiting for them to converge.
-Figure 5 shows significant voltage limit violations without control and demonstrates that our proposed distributed nested feedback optimization approach successfully enforces the voltage limit even for the most remote, and therefore, the most sensitive bus.
-An average voltage violation (AVV), defined in Fig.
-5, of $2.1\times 10^{-4}$ pu is achieved, which is only higher than that of the centralized approach.
+IV-C Dynamic case results
+In the dynamic case, we study the performance of the approaches in a time-varying environment, where the approaches are implemented online to track the time-varying optimizers of (6) without waiting for them to converge.
+Figure 8a shows significant voltage limit violations without control.
+Figures 8b and 8c demonstrate that the centralized and our proposed distributed nested feedback optimization approaches both successfully enforce the voltage limit, although temporary voltage limit violations are inevitable due to the corrective nature of these algorithms.
+The simple two-metric approach and the heuristic approach in [23] result in prolonged voltage limit violations as seen in Figs.
+8d and 8e, respectively.
+Figure 8f shows that the distributed approach in [24] does not track the time-varying optimizers well and experiences oscillations at the end of the simulation period.
+The discrepancy between the actual setpoints and algorithm iterates (without projection) potentially renders the algorithm slow and unstable.
+Figure 8g shows that the local droop control approach does not enforce the voltage limit due to lack of coordination.
+Finally, for clarity, Fig.
+8h presents voltages for the most remote, and therefore, the most sensitive bus (bus 95) under different approaches.
+Furthermore, Fig.
+9 shows reactive power profiles under different approaches.
+Shown in Figs.
+9a and 9b, our proposed distributed approach achieves similar profiles with the centralized one, with an average setpoint deviation of $-1.3\times 10^{-3}$ kVar.
+The two-metric approach and the heuristic approach do not work well in this case study.
+The approach in [24] experiences oscillations.
+The local droop control approach does not in theory or practice track the optimizers of (6).
+Finally, we define the average voltage violation (AVV) in (20), where $k$ is the iteration index and $K$ is the total number of iterations.
+$\operatorname{AVV}=\frac{1}{K}\sum_{k=1}^{K}\left([\tilde{v}^{k}-\overline{v}]^{+}+[\underline{v}-\tilde{v}^{k}]^{+}\right).$ (20)
+TABLE I: Average voltage violations for various approaches.
+Approach AVV (pu) w.r.t. centralized
+No control $1.5\times 10^{-2}$ 192x
+Centralized $8.0\times 10^{-5}$ 1x
+Proposed $1.6\times 10^{-4}$ 2x
+Two-metric $2.5\times 10^{-3}$ 31x
+Tang 2024 [23] $3.3\times 10^{-3}$ 41x
+Qu 2020 [24] $8.6\times 10^{-4}$ 11x
+Droop control [37, 38] $4.5\times 10^{-4}$ 6x
+Table I shows the AVV values for bus 95 (the most sensitive bus) under various approaches.
+An AVV of $1.6\times 10^{-4}$ pu is achieved for the proposed approach, which is only higher than that of the centralized approach.
 The latter however requires a centralized communication architecture and is thus susceptible to single-point failures.
-The proposed nested feedback approach significantly outperforms other distributed ones, which is visible in Fig.
-5 and quantified using the AVV.
-IV Conclusion
+The proposed nested feedback approach significantly outperforms other existing distributed or local approaches.
+V Conclusion
 In this paper, we focused on the distributed voltage regulation problem in distribution systems and proposed a nested feedback optimization approach.
 We theoretically analyzed its convergence.
-Our simulation results showed that the approach, while only requiring short-range communication between physical neighbours, achieved satisfactory voltage regulation and outperformed other distributed approaches.
+Our simulation results showed that the approach, while only requiring short-range communication between physical neighbours, achieved satisfactory voltage regulation and outperformed existing distributed and local approaches.
 The proposed distributed approach does not extend to congestion management while the centralized does so.
@@
 Finally, the approach works for either active or reactive power control, but not joint control.
-This is because the observation in Section II-D does not hold anymore since the voltage difference is no longer the gradient of either active or reactive power injection, but represents their joint non-separable impact.
+This is because the observation in Section III-B does not hold anymore since the voltage difference is no longer the gradient of either active or reactive power injection, but represents their joint non-separable impact.
 However, one might deploy active and reactive power feedback controllers separately at different sampling time.

### v4

@@
-Distributed Online Feedback Optimization for Real-time Distribution System Voltage Regulation
+Link to IEEEXplore: https://ieeexplore.ieee.org/document/11066276
 Abstract
@@
 This paper proposes a distributed OFO design based on a nested feedback optimization strategy and analyzes its convergence.
+The strategy preserves end-users’ privacy by keeping voltage data local.
 Numerical study results demonstrate that the proposed design achieves effective voltage regulation and outperforms other distributed and local approaches.
@@
 In contrast to deep reinforcement learning [13, 14], OFO does not need a complicated (often centralized offline) training process while offering theoretical guarantees to obtain globally optimal solutions.
+Recent studies in [15, 16, 17, 18] have explored data-driven approaches to train linear and nonlinear local voltage controllers via deep learning, reinforcement learning, and unsupervised learning.
+However, these approaches demand extensive training data and lack optimality guarantees without real-time coordination between DERs.
 Furthermore, unlike using traditional voltage regulators such as on-load tap changers and capacitor banks [19], OFO provides rapid voltage regulation with inverter-interfaced DERs and can complement these slow-reacting devices in real time [20].
-While many OFO controllers rely on centralized communication [1, 17, 2, 8, 4, 5, 18, 9, 19, 20, 21], distributed OFO based on communication between physical neighbours appears to be a compelling approach due to its robustness to single-point failures[6, 3, 16, 22, 23, 24].
+While many OFO controllers rely on centralized communication [1, 21, 2, 8, 4, 5, 22, 9, 23, 24, 25], distributed OFO [6, 3, 20, 26, 27, 28] based on communication between physical neighbours appears to be a compelling approach due to its robustness to single-point failures and enhancement of privacy.
 Among the existing distributed designs, [6, 3, 20] demand particular forms of objective functions, while [26] relies on disseminating global information through neighbouring nodes.
@@
 This heuristic approach does not assure adequate voltage regulation.
-In this paper, we propose a distributed OFO design inspired by [24], where the distributed implementation was enabled by the sparsity of the inverse of the network sensitivity matrix $\mathbf{X}$, i.e.
+The motivation of this study is to develop a real-time voltage regulation strategy that optimally coordinates DERs on a fast timescale and operates without requiring an accurate grid model or load data.
+Furthermore, the strategy should be robust to single-point failures, meaning that unlike centralized approaches—where the failure of a single coordination unit can disrupt the entire system—the approach should ensure that voltage regulation remains functional even if some nodes fail.
+Finally, the strategy should preserve privacy.
+Exposing nodal voltage data could allow external entities to infer sensitive load information, which is typically considered private.
+In this paper, we propose a distributed OFO design that meets these requirements.
+The strategy is inspired by [28], where the distributed implementation was enabled by the sparsity of the inverse of the network sensitivity matrix $\mathbf{X}$, i.e.
 $\mathbf{X}^{-1}$ is sparse.
@@
 Nevertheless, this strategy takes a prohibitively large number of iterations to converge, which may jeopardize its performance in online implementation.
-We tackle this challenge by proposing an iterative algorithm to solve this non-Euclidean projection problem that requires only local information and operates in an online feedback-based manner.
-This leads to our overall nested feedback optimization approach, where the outer loop comprises the OFO iterations requiring communication between physical neighbours and yielding tentative but not necessarily feasible DER setpoints, while the inner loop solves the non-Euclidean projection problem to map the tentative setpoints to actual feasible setpoints using local information.
+We tackle this challenge by proposing an iterative projected gradient descent algorithm to solve this non-Euclidean projection problem where each node only requires information from its neighbours.
+This leads to our overall nested feedback optimization approach, where the outer loop comprises the OFO iterations requiring communication between physical neighbours and yielding tentative but not necessarily feasible DER setpoints, while the inner loop solves the non-Euclidean projection problem to map the tentative setpoints to actual feasible setpoints.
+Unlike the centralized approaches, our method is robust to single-point failures as it operates without a central coordination unit and preserves privacy.
+Compared to [3, 6, 20], our approach accommodates a broader range of objective functions.
+In contrast to [26, 27], our approach keeps voltage information local enhancing privacy.
+Finally, compared to [28], our approach converges in significantly fewer iterations and achieves superior voltage regulation performance.
 To summarize, the main contributions of this paper are:
-  • A local iterative algorithm is proposed to solve the non-Euclidean projection problem that operates in an online feedback-based manner.
-  • An efficient nested feedback optimization strategy that ensures adequate voltage regulation while eliminating the need for an accurate grid model and consumption of non-controllable loads is developed.
-  • The proposed strategy is theoretically analyzed and numerically assessed on a 96-bus system with high-resolution generation and load data.
+  • We develop a projected gradient descent algorithm that exchanges information between neighboring nodes in a backward-forward manner to solve the non-Euclidean projection problem.
+  • We propose a nested feedback optimization strategy for distribution system voltage regulation that operates without requiring an accurate grid model or load data.
+The strategy optimally coordinates DERs on a fast timescale while ensuring robustness to single-point failures and preserving privacy by keeping voltage data local.
+  • We theoretically prove the convergence of the proposed strategy.
+Simulation results demonstrate that it achieves comparable voltage regulation to its centralized counterpart while significantly outperforming other distributed and local approaches.
 The remainder of this paper is structured as follows: Section II presents the system modeling, problem formulation, and a centralized OFO strategy.
@@
 The $N\times N$-dimensional symmetrical matrices $\mathbf{R}$ and $\mathbf{X}$ are given in (3), where $\mathcal{E}_{i}$ represents the set of cables on the unique path from the substation bus to bus $i$.
-$\mathbf{R}_{ij}=\sum_{(h,k)\in\mathcal{E}_{i}\cap\mathcal{E}_{j}}r_{hk},\mathbf{X}_{ij}=\sum_{(h,k)\in\mathcal{E}_{i}\cap\mathcal{E}_{j}}x_{hk}.$ (3)
+$R_{ij}=\sum_{(h,k)\in\mathcal{E}_{i}\cap\mathcal{E}_{j}}r_{hk},X_{ij}=\sum_{(h,k)\in\mathcal{E}_{i}\cap\mathcal{E}_{j}}x_{hk}.$ (3)
 These matrices accordingly capture the network sensitivities, i.e.
@@
 Furthermore, [28] shows that their inverse matrices $\mathbf{R}^{-1}$ and $\mathbf{X}^{-1}$ have the sparsity pattern such that
-$\mathbf{R}_{ij}^{-1}\neq 0,\mathbf{X}_{ij}^{-1}\neq 0\iff(i,j)\in\mathcal{E}\text{ or }i=j.$ (5)
+$R_{ij}^{-1}\neq 0,X_{ij}^{-1}\neq 0\iff(i,j)\in\mathcal{E}\text{or }i=j.$ (5)
 Figure 2 visualizes $\mathbf{X}$ and $\mathbf{X}^{-1}$ for the test system in Section IV, where the sparsity patterns are shown.
@@
 $\underline{q}_{i}\leq q_{i}\leq\overline{q}_{i},\forall i\in\mathcal{N}^{+}.$ (6c)
+Remark 1.
+The objective function is designed to enable our distributed voltage regulation algorithm.
+Since the algorithm relies solely on local information and data from neighboring nodes, the gradient of the objective function must be locally computable.
+This requirement excludes certain objectives, such as loss minimization, which would require a centralized communication structure.
+Additionally, in case of overvoltage due to PV generation, minimizing reactive power usage contributes to loss reduction.
+When PV inverters absorb reactive power to lower voltage magnitudes, they increase reactive power transfer within the system, which already experiences reactive power consumption from non-controllable loads and cable losses.
+The more reactive power PV inverters absorb, the higher the system losses.
+Thus, minimizing their reactive power consumption helps reduce overall losses, as demonstrated by numerical results in Section IV-C.
+Finally, this objective function also enhances inverter reliability.
+As noted in [33], operating inverters at power factors significantly different from unity can reduce their lifespan.
+Therefore, our approach aligns with the interest of end-users by promoting longer inverter lifetime.
 For further developments, the dual problem of (6) is formulated in (7).
@@
 Thus, this section introduces a feedback controller based on OFO.
-The controller leverages primal-dual gradient projection (PDGP) as the underlying algorithm [1, 17, 2] to track the time-varying optimizers of (6), which requires a centralized gather-and-broadcast communication architecture.
+The controller leverages primal-dual gradient projection (PDGP) as the underlying algorithm [1, 21, 2] to track the time-varying optimizer of (6), which requires a centralized gather-and-broadcast communication architecture.
+Compared to other dual optimization approaches such as dual ascent and the alternating method of multipliers (ADMM), PDGP offers additional tuning of primal step sizes, which is beneficial under fluctuating grid conditions, and enables faster implementation of DER setpoints.
+The centralized OFO-based solution is included primarily as a baseline for comparison and as a foundation for the distributed methods developed later.
 Specifically, the PDGP algorithm performs projected gradient ascent and descent steps iteratively for the dual and primal variables, respectively.
@@
     $q_{i}^{k+1}=\big{[}q_{i}^{k}-\alpha[c_{i}q_{i}^{k}+[\mathbf{X}(\boldsymbol{\lambda}^{k+1}-\boldsymbol{\mu}^{k+1}+r^{p}\mathbf{q}^{k})]_{i}]\big{]}_{\underline{q}_{i}^{k}}^{\overline{q}_{i}^{k}}.$ (9c)
+Remark 2.
+The regularization ensures strong convexity of the problem and guarantees convergence.
+However, it introduces a deviation from the exact solution of (7), resulting in a close approximation instead.
+As shown in [34], the discrepancy is bounded and proportional to $\sqrt{r_{d}}$.
+A practical guide for choosing the regularization parameters is to start with small values and gradually increase them until the algorithm stabilizes.
 III Distributed OFO
@@
 However, this approach requires an excessively large number of iterations to reach convergence, significantly hindering its performance in online implementation.
-Figure 4: Illustration of the approximation.
-To tackle this challenge, this paper proposes an iterative algorithm to solve (13) that requires only local information and is implemented in an online feedback-based manner.
+To tackle this challenge, this paper proposes an iterative projected gradient descent (PGD) algorithm to solve (13) that exchanges information only between physical neighbours in a backward-forward manner.
 This leads to our nested feedback optimization approach, where the outer loop comprises the OFO iterations giving tentative reactive power setpoints, and the inner loop solves the non-Euclidean projection problem (13) to generate feasible setpoints.
-Our key observation is that given a feasible $\mathbf{u}^{\tau}\in\mathcal{X}^{k}$, the gradient of the objective function in (13) at $\mathbf{u}^{\tau}$ is $\mathbf{X}(\mathbf{u}^{\tau}-\mathbf{\dot{q}}^{k+1})$, which can be approximated by the voltage difference $\mathbf{v}(\mathbf{u}^{\tau})-\mathbf{v}(\mathbf{\dot{q}}^{k+1})$ according to (2) assuming system disturbances do not change significantly when all nodes are locally implementing this inner iterative algorithm, i.e.
-$\mathbf{X}(\mathbf{u}^{\tau}-\mathbf{\dot{q}}^{k+1})\approx\mathbf{v}(\mathbf{u}^{\tau})-\mathbf{v}(\mathbf{\dot{q}}^{k+1}).$ (14)
-The proposed algorithm is based on the non-scaled Euclidean gradient projection and proceeds as:
-  1. 1.
-    Implement
-    $\mathbf{q}^{k}_{\epsilon}=\mathbf{q}^{k}+\epsilon(\mathbf{\dot{q}}^{k+1}-\mathbf{q}^{k}),$ (15)
-    and measure the voltage $\mathbf{\tilde{v}}(\mathbf{q}^{k}_{\epsilon})$. Estimate
-    $\mathbf{\tilde{v}}(\mathbf{\dot{q}}^{k+1})\approx\mathbf{\tilde{v}}^{k}+(\mathbf{\tilde{v}}(\mathbf{q}^{k}_{\epsilon})-\mathbf{\tilde{v}}^{k})/\epsilon.$ (16)
-  2. 2.
-    Given a feasible $\mathbf{u}^{\tau}$, implement it, measure the voltage $\mathbf{\tilde{v}}(\mathbf{u}^{\tau})$, compute the next implementable reactive power setpoint $\mathbf{u}^{\tau+1}$ as (17), and repeat this step for $T$ iterations.
-    $\mathbf{u}^{\tau+1}=[\mathbf{u}^{\tau}-\alpha^{u}(\mathbf{\tilde{v}}(\mathbf{u}^{\tau})-\mathbf{\tilde{v}}(\mathbf{\dot{q}}^{k+1}))]_{\mathbf{\underline{q}}^{k}}^{\mathbf{\overline{q}}^{k}}.$ (17)
-Remark 1.
-In the first step, $\epsilon$ represents a small exploration parameter, which is similar to that in model-free feedback optimization [29, 30, 31, 32].
-Since $\mathbf{\dot{q}}^{k+1}$ can be pretty far away from the feasible set, it might be impossible to implement it.
-Therefore, we implement $\mathbf{q}^{k}_{\epsilon}$ and estimate $\mathbf{\tilde{v}}(\mathbf{\dot{q}}^{k+1})$ based on it.
-This is illustrated in Fig.
-4.
-This could still result in minor temporary DER capacity limit violations, which should be manageable by the DERs.
-To ensure strict capacity limit satisfaction, the original feasible set can be slightly deflated [32].
-The initial feasible $\mathbf{u}^{\tau}$ in the second step can be chosen as $\mathbf{q}^{k}$.
-The online implementation of the overall nested algorithm is summarized in Algorithm 1.
-Data: $\mathbf{q}^{0}$
-Result: $\mathbf{q}^{k}$, $k\geq 0$; $\mathbf{u}^{\tau}$, $\tau\leq T$
-1 Set k = 0;
+Given a $\mathbf{u}^{k,\tau}\in\mathcal{X}^{k}$ where $\tau=0,1,2,\cdots,T$ is the index for the inner loop, the PGD algorithm proceeds to compute $\mathbf{u}^{k,\tau+1}\in\mathcal{X}^{k}$ using (14) where $\mathbf{X}(\mathbf{u}^{k,\tau}-\mathbf{\dot{q}}^{k+1})$ is the gradient of the objective function of (13) at $\mathbf{u}^{k,\tau}$.
+$\mathbf{u}^{k,\tau+1}=[\mathbf{u}^{k,\tau}-\alpha^{u}\mathbf{X}(\mathbf{u}^{k,\tau}-\mathbf{\dot{q}}^{k+1})]_{\mathbf{\underline{q}}^{k}}^{\mathbf{\overline{q}}^{k}}.$ (14)
+The remaining challenge is to compute the gradient in a distributed manner.
+To achieve this, we exploit the underlying tree topology of radial distribution networks, which enables information propogation in a backward-forward manner.
+In [35], this procedure is derived using LDU factorization.
+While backward-forward strategies have been used on tree graphs, our contribution lies in leveraging them to solve the non-Euclidean projection problem, completing a distributed voltage control scheme that optimally coordinates DERs on a fast timescale, operates in a feedback-based manner, and preserves end-users’ privacy by keeping voltage data local.
+For notational simplicity, we denote $\mathbf{u}^{k,\tau}-\mathbf{\dot{q}}^{k+1}$ as $\boldsymbol{\omega}^{k,\tau}$ and denote its $i$-th element as $\omega_{i}^{k,\tau}$, which is locally known by node $i$.
+In the backward step, for each node $i$, upon collecting $\xi_{j}^{k,\tau}$ for all $j\in\mathcal{C}_{i}$ where $\mathcal{C}_{i}$ represents the set of child nodes of $i$, it sends to its unique parent node
+$\xi_{i}^{k,\tau}=\sum_{j\in\mathcal{C}_{i}}\xi_{j}^{k,\tau}+\omega_{i}^{k,\tau}.$ (15)
+Note that this step runs recursively from the farthest nodes up to but not including the closest nodes to the point of common coupling (PCC), where $\xi_{j}^{k,\tau}$ is recursively defined.
+In a nutshell, $\xi_{i}^{k,\tau}$ represents the sum of $\omega_{n}^{k,\tau}$ where $n$ belongs to all of its downstream nodes including itself.
+In the forward step, for each node $i$, upon receiving $\zeta_{h}^{k,\tau}$ from its unique parent node, it sends to its child node(s)
+$\zeta_{i}^{k,\tau}=\zeta_{h}^{k,\tau}+x_{hi}\xi_{i}^{k,\tau}.$ (16)
+Similarly, this step runs recursively from the closest nodes to the PCC up to but not including the farthest nodes, where $\zeta_{h}^{k,\tau}$ is recursively defined.
+Likewise, $\zeta_{i}^{k,\tau}$ represents the sum of $x_{mn}\xi_{n}^{k,\tau}$ where $n$ belongs to all of its upstream nodes including itself.
+Based on the information received from its child node(s) and parent node, each node $i$ can then determines its entry in $\mathbf{X}\boldsymbol{\omega}^{k,\tau}$ as
+$[\mathbf{X}\boldsymbol{\omega}^{k,\tau}]_{i}=\zeta_{h}^{k,\tau}+x_{hi}\xi_{i}^{k,\tau}.$ (17)
+Note that in this process, each node only communicates with its physical neighbours once. The overall nested feedback optimization algorithm is summarized in Algorithm 1.
+Data: $\mathbf{q}^{0}$, $\boldsymbol{\lambda}^{0}$, $\boldsymbol{\mu}^{0}$
+Result: $\mathbf{q}^{k}$, $k\geq 0$
+1Set k = 0;
 2 while $k\geq 0$ do
 3       Implement $\mathbf{q}^{k}$ and collect voltage measurement $\tilde{\mathbf{v}}(\mathbf{q}^{k})$;
+4       Update $\boldsymbol{\lambda}^{k+1}$ and $\boldsymbol{\mu}^{k+1}$ via (9a)-(9b);
 5       Communicate $\mathbf{q}^{k}$ between physical neighbours and calculate $\dot{\mathbf{q}}^{k+1}$ via (11);
-5       Implement $\mathbf{q}^{k}_{\epsilon}$, measure voltage $\mathbf{\tilde{v}}(\mathbf{q}^{k}_{\epsilon})$ and estimate $\mathbf{\tilde{v}}(\mathbf{\dot{q}}^{k+1})$;
 6       Set $\tau=0$;
 7       while $\tau<T$ do
-8             Implement $\mathbf{u}^{\tau}$ and measure voltage $\tilde{\mathbf{v}}(\mathbf{u}^{\tau})$ (choose $\mathbf{u}^{0}$ as $\mathbf{q}^{k}$);
-9             Calculate next implementable $\mathbf{u}^{\tau+1}$ via (17);
+8            Use the backward-forward algorithm (15)-(17) to compute $\mathbf{X}(\mathbf{u}^{k,\tau}-\mathbf{\dot{q}}^{k+1})$ (choose $\mathbf{u}^{k,0}$ as $\mathbf{q}^{k}$);
+9             Calculate next $\mathbf{u}^{\tau+1}$ via (14);
 10             Set $\tau=\tau+1$;
@@
 Algorithm 1 Nested feedback optimization algorithm
+Remark 3.
+In principle, the same approach could be applied to (9c) to develop a distributed strategy.
+However, this would require sharing dual variables.
+Due to the correlation between voltage and active power, exposing these variables could enable external entities to infer sensitive load information, which is typically considered private.
+In contrast, our strategy avoids sharing dual variables.
+Remark 4.
+The proposed method assumes that each DER has access to topological information, such as knowledge of its immediate neighbors.
+This information is assumed to be obtained during the algorithm’s planning phase.
+However, we acknowledge that this may not always be feasible.
+Therefore, exploring approaches that reduce or eliminate this dependency represents an interesting research direction to enhance the practicality and adaptability of the algorithm.
+Remark 5.
+The addition or removal of DERs may require reconfiguration of the communication graph between agents.
+However, if each node is interpreted as a virtual DER—potentially representing an aggregation point in the network—the communication structure can remain unchanged.
+Under this setting, the presence or absence of individual physical DERs does not necessarily impact the agent-level communication.
 III-C Convergence analysis
@@
 The linear relation (2) holds between voltages and reactive power injections.
-Based on the first assumption, $\mathbf{v}(\mathbf{u}^{\tau})-\mathbf{v}(\mathbf{\dot{q}}^{k+1})$ is the exact gradient of the objective function in (13) at $\mathbf{u}^{\tau}$.
-The inner loop then represents a standard Euclidean gradient projection algorithm to solve the quadratic program (13).
-Consequently, $\mathbf{u}^{\tau}$} is a converging sequence by invoking Proposition 2.3.2 in [25] with $0<\alpha^{u}<2/\lambda_{max}(\mathbf{X})$ where $\lambda_{max}(\mathbf{X})$ is the largest eigenvalue of $\mathbf{X}$.
 Assumption 2.
 The inner iterations reach convergence after $T$ iterations.
-The overall nested algorithm thus represents a scaled PDGP algorithm to solve (6).
+The inner loop represents a standard Euclidean gradient projection algorithm to solve the quadratic program (13).
+Note that the backward-forward step is only used to gather gradient information; each agent then performs a projected gradient descent step.
+Consequently, $\mathbf{u}^{\tau}$} is a converging sequence by invoking Proposition 2.3.2 in [29] with $0<\alpha^{u}<2/\lambda_{max}(\mathbf{X})$ where $\lambda_{max}(\mathbf{X})$ is the largest eigenvalue of $\mathbf{X}$.
+Since we cannot choose $T$ to be arbitrarily large in a practical system, this algorithm will generate an approximate solution to (13).
+When $T$ is chosen sufficiently large, the algorithm will converge arbitrarily close to the optimum of (13) and the approximation error will be negligible.
+In Section IV, we demonstrate its convergence with numerical results.
+Under the two assumptions, the overall nested algorithm represents a scaled PDGP algorithm to solve (6).
 By a transformation of variables defined by: $\mathbf{q}^{\prime}=\mathbf{X}^{\frac{1}{2}}\mathbf{q}$ and $\mathbf{v}^{\prime}=\gamma\mathbf{v}$ where $\gamma=\sqrt{\frac{\alpha^{d}}{\alpha}}$ is the squared root of the ratio between the dual and primal step sizes, the problem (6) can be cast as an equivalent quadratic program in the space of $\mathbf{q}^{\prime}$ and $\mathbf{v}^{\prime}$ as in (18).
@@
 $\qquad\cdot\mathbf{C}\mathbf{X}^{-\frac{1}{2}}\mathbf{q^{\prime}}^{k}+\gamma\mathbf{X}^{\frac{1}{2}}(\boldsymbol{\lambda^{\prime}}^{k+1}-\boldsymbol{\mu^{\prime}}^{k+1})+r^{p}\mathbf{q^{\prime}}^{k}]\big{]}\rVert_{2}^{2}.$
-which are equivalent to (9a)-(9b) and (11)-(13) [25, Eqs.
+which are equivalent to (9a)-(9b) and (11)-(13) [29, Eqs.
 (2.37) and (2.38)].
@@
 Remark 6.
-First, while the linear voltage-reactive power relation is assumed to synthesize the controller and analyze its convergence, AC power flow is run in the simulation to calculate the actual voltages.
+While the linear voltage-reactive power relation is assumed to synthesize the controller and analyze its convergence, AC power flow is run in the simulation to calculate the actual voltages.
 Such an assumption was also made in the proofs of [3, 27].
-Second, a large $T$ is not necessary in the simulation.
-We use $T=4$ so the inner loop might not converge.
-However, the algorithm performs well in the simulation.
-Notably, a similar assumption was made in [3].
 IV Case study
@@
 Figure 4: A 96-bus test system from Simbench [36].
+(a) Trajectories of the cost function
+(b) Distribution of the relative cost function (w.r.t. the optimum)
+Figure 5: Results of 1000 random problem instances.
 In the numerical simulation, we study a 100% penetration scenario of photovoltaics (PVs) in a 96-bus system, i.e.
@@
 The test system is adopted from Simbench [36] and is shown in Fig.
-5, which is expected to experience overvoltage issues.
+4, which is expected to experience overvoltage issues.
 The network topology and impedance data are the same as those in the original dataset.
@@
 The unit for power is kW/kVar.
-The step sizes are chosen with a trial-and-error strategy [24] as $\alpha^{d}=10^{6}$, $\alpha=5\times 10^{-4}$, and $\alpha^{u}=10^{2}$.
-The exploration parameter $\epsilon$ is chosen as $10^{-5}$, resulting in a maximum DER reactive power limit violation of 0.6%.
-The primal and dual regularization is used to establish the theoretical convergence result.
-The impact of some small regularization parameters on numerical results is negligible.
-In the simulation, we set them as $10^{-4}$.
-For each generation and load data point, an OFO iteration is run which yields tentative reactive power setpoints, followed by one exploration step and 4 inner feedback iterations of (17), i.e.
-we assume that every setpoint in Algorithm 1 is implemented for 1 second.
+For the inner loop, $\alpha^{u}$ is chosen as $0.99\cdot 2/\lambda_{max}(\mathbf{X})$.
+For the outer loop, the step sizes are chosen with a trial-and-error strategy [28] as $\alpha^{d}=10^{6}$ and $\alpha=3\times 10^{-4}$.
+Following the guideline in Remark 2, the regularization parameters $r_{p}$ and $r_{d}$ are set to $10^{-4}$.
+The algorithm performs well with such small regularization, likely due to the strongly convex objective function in (6).
+For each generation and load data point, 3 OFO iterations are run which yields tentative reactive power setpoints, followed by 10 inner iterations, i.e.
+we assume that each inner iteration takes 200 milliseconds.
+It is crucial that the system responds faster than the DER setpoint update rate.
+In our application, setpoints are updated on a (sub-)second timescale, and the system is assumed to react even more quickly.
+Similarly, [40, 1, 2] consider DER update time of 0.1, 0.33, and 1 second, respectively.
+Another relevant example is droop control, where voltage measurements and setpoint updates occur in real time, typically within the subsecond range [41].
 Finally, the nested algorithm requires only basic arithmetic operations and is computationally very efficient.
 For a series implementation of a 4-hour simulation, i.e.
-in total 14400 iterations, it takes 8.1 seconds, averaging 0.56 milliseconds per iteration.
+in total 7200 iterations, it takes 6.7 seconds, averaging 0.93 milliseconds per iteration.
 Figure 6: Voltage profiles over 500 iterations for several buses in the static case using the centralized and the proposed approaches.
@@
 Figure 7: Reactive power profiles after 500 iterations for the centralized and the proposed approaches.
+Figure 8: Loss comparison between the objectives of reactive power minimization and loss minimization.
 (a) No control
@@
 Figure 10: Reactive power profiles under different approaches in the dynamic simulation.
-IV-B Static case results
+IV-B Convergence of inner loop
+To demonstrate the convergence of the inner loop, 1000 random problem instances for (13) are generated.
+Figure 5a illustrates that the cost function drops significantly within a few iterations, leading to rapid algorithm convergence.
+Additionally, Fig.
+5b depicts the distribution of the relative cost function with respect to the optimum.
+After approximately 10 iterations, the cost function stabilizes.
+Thus, in this study, we set the inner loop to run for 10 iterations, ensuring a balance between solution accuracy of the inner loop and the update frequency of the outer loop.
+IV-C Static case results
 In the static case, the generation and load data are kept unchanged to study the convergence of the nested approach.
 The simulation uses data at 12:00.
-Figure 6 shows that the proposed approach suppresses the initial voltage limit violations quickly and converges after around 200 total iterations, i.e.
-33 outer OFO iterations.
+Figure 6 shows that the proposed approach suppresses the initial voltage limit violations quickly and converges after around 60 outer iterations.
 The approach also utilizes the available voltage limit as efficiently as the centralized one.
@@
 This is reasonable since the gradient scaling only alters the converging trajectory, but does not change the solution to (6).
-IV-C Dynamic case results
-In the dynamic case, we study the performance of the approaches in a time-varying environment, where the approaches are implemented online to track the time-varying optimizers of (6) without waiting for them to converge.
+Finally, Fig.
+8 demonstrates that under overvoltage issues caused by excessive PV generation, the chosen objective function of minimizing reactive power usage aligns with minimizing system losses.
+However, under undervoltage conditions, the outcome may differ, as increased reactive power injection could potentially help lower system losses.
+This highlights a tradeoff between the distributed communication architecture and the flexibility in selecting optimization objectives.
+IV-D Dynamic case results
+In the dynamic case, we study the performance of the approaches in a time-varying environment, where the approaches are implemented online to track the time-varying optimizer of (6) without waiting for them to converge.
 Figure 9a shows significant voltage limit violations without control.
-Figures 8b and 8c demonstrate that the centralized and our proposed distributed nested feedback optimization approaches both successfully enforce the voltage limit, although temporary voltage limit violations are inevitable due to the corrective nature of these algorithms.
+Figures 9b and 9c demonstrate that the centralized and our proposed distributed nested feedback optimization approaches both successfully enforce the voltage limit, although temporary voltage limit violations are inevitable due to the corrective nature of these algorithms.
 The simple two-metric approach and the heuristic approach in [27] result in prolonged voltage limit violations as seen in Figs.
-8d and 8e, respectively.
+9d and 9e, respectively.
 Figure 9f shows that the distributed approach in [28] does not track the time-varying optimizers well and experiences oscillations at the end of the simulation period.
-The discrepancy between the actual setpoints and algorithm iterates (without projection) potentially renders the algorithm slow and unstable.
+The discrepancy between the actual setpoints and the algorithm iterates (without projection) potentially renders the algorithm slow and unstable.
 Figure 9g shows that the local droop control approach does not enforce the voltage limit due to lack of coordination.
 Finally, for clarity, Fig.
-8h presents voltages for the most remote, and therefore, the most sensitive bus (bus 95) under different approaches.
+9h presents voltages for the most remote, and therefore, the most sensitive bus (bus 95) under different approaches.
 Furthermore, Fig.
@@
 Shown in Figs.
-9a and 9b, our proposed distributed approach achieves similar profiles with the centralized one, with an average setpoint deviation of $-1.3\times 10^{-3}$ kVar.
+10a and 10b, our proposed distributed approach achieves similar profiles with the centralized one, with an average setpoint deviation of 0.01 kVar.
 The two-metric approach and the heuristic approach do not work well in this case study.
 The approach in [28] experiences oscillations.
-The local droop control approach does not in theory or practice track the optimizers of (6).
+The local droop control approach does not in theory or in practice track the optimizer of (6).
 Finally, we define the average voltage violation (AVV) in (20), where $k$ is the iteration index and $K$ is the total number of iterations.
 $\operatorname{AVV}=\frac{1}{K}\sum_{k=1}^{K}\left([\tilde{v}^{k}-\overline{v}]^{+}+[\underline{v}-\tilde{v}^{k}]^{+}\right).$ (20)
-TABLE I: Average voltage violations for various approaches.
-Approach AVV (pu) w.r.t. centralized
-No control $1.5\times 10^{-2}$ 192x
-Centralized $8.0\times 10^{-5}$ 1x
-Proposed $1.6\times 10^{-4}$ 2x
-Two-metric $2.5\times 10^{-3}$ 31x
-Tang 2024 [23] $3.3\times 10^{-3}$ 41x
-Qu 2020 [24] $8.6\times 10^{-4}$ 11x
-Droop control [37, 38] $4.5\times 10^{-4}$ 6x
-Table I shows the AVV values for bus 95 (the most sensitive bus) under various approaches.
-An AVV of $1.6\times 10^{-4}$ pu is achieved for the proposed approach, which is only higher than that of the centralized approach.
-The latter however requires a centralized communication architecture and is thus susceptible to single-point failures.
+TABLE I: Average voltage violations and losses for various approaches.
+Approach AVV (pu) AVV w.r.t. central. Loss (kWh)
+No control $1.5\times 10^{-2}$ 192x 48.2
+Centralized $8.0\times 10^{-5}$ 1x 64.0
+Proposed $9.9\times 10^{-5}$ 1.2x 64.0
+Two-metric $2.5\times 10^{-3}$ 31x 59.3
+Tang 2024 [27] $3.3\times 10^{-3}$ 41x 57.4
+Qu 2020 [28] $8.6\times 10^{-4}$ 11x 65.9
+Droop control [42] $4.5\times 10^{-4}$ 6x 67.8
+Table I shows the AVV values for bus 95 (the most sensitive bus) and system losses under various approaches.
+An AVV of $9.9\times 10^{-5}$ pu is achieved for the proposed approach, which is slightly higher than that of the centralized approach.
+The latter however requires a centralized communication architecture and is thus susceptible to failure of the central coordination unit.
 The proposed nested feedback approach significantly outperforms other existing distributed or local approaches.
+Because reactive power is absorbed to mitigate voltage rise, system losses increase.
+The proposed approach results in the same losses as the centralized method.
+IV-E Impact of measurement noise
+Figure 11: Voltages for bus 95 under measurement noise.
+This section examines the impact of measurement noise on the voltage regulation performance.
+The noise is modeled using a zero-mean Gaussian distribution with a standard deviation of $\sigma$.
+Figure 11 illustrates its effect.
+As measurement noise increases, voltage fluctuations become more pronounced.
+Nevertheless, the controller successfully maintains voltages within safe limits.
+The temporary voltage limit violations can be mitigated by setting a tighter limit, e.g.
+1.048 pu.
+IV-F Impact of node failures
+Figure 12: Voltages for bus 95 under a failed node.
+Figure 13: Impact of node failres on AVV.
+In this section, we demonstrate the robustness of the proposed strategy against random node failures.
+When a node fails, its neighbouring nodes are assumed to bypass it and continue communication.
+Figure 12 illustrates an example voltage profile with a failed node, showing that the system maintains the voltage limit despite a random node failure.
+Figure 13 depicts how AVV increases with the number of failed nodes but remains within a reasonable range, indicating the algorithm’s robustness.
+Unlike centralized approaches, where failure of the central coordination unit leads to a complete loss of voltage regulation, the proposed method continues to function effectively.
+Enhancing system resilience through the development of failure detection and bypass mechanisms is an important direction for future research.
+IV-G Impact of update frequency
+Figure 14: AVV under various update frequency.
+Figure 14 quantifies the impact of the OFO controller update frequency on its voltage regulation performance.
+Note that the simulated generation and load data have a resolution of 6 seconds.
+As expected, more frequent updates improve performance.
+However, even with an outer iteration length of 24 seconds, the performance deterioration remains limited.
+This suggests that the method retains effectiveness even under slower update rates.
+While we assume relatively fast update in the inner loop (e.g.
+200 milliseconds per iteration), we acknowledge that this may be optimistic in large or deep networks.
+Therefore, this section serves to assess the controller’s robustness to slower updates, which relaxes the need for low-latency communication.
+The optimal update frequency depends on the communication infrastructure and requires a cost-effectiveness analysis.
+Future work will focus on analyzing the scalability of the controller under more realistic communication constraints, including increased latencies and asynchronous operation.
+Figure 15: Voltages for bus 95 under various objective functions.
+IV-H Impact of the objective function
+In this section, we compare our objective function with the alternative objective of minimizing $\mathbf{q}^{\intercal}\mathbf{X}\mathbf{q}$, as adopted in [3, 12, 20].
+Figure 15 shows that this weighted objective function (i.e.
+weighted by $\mathbf{X}$) also effectively enforces the voltage limit in the dynamic case.
+The voltage profile obtained using the weighted cost function yields a lower average voltage violation (AVV) of $3.0\times 10^{-5}$ pu but exhibits greater variation.
+Given that the system faces overvoltage issues, the voltage constraints are expected to be binding at the optimal solution.
+Accordingly, we use 1.05 pu as the reference voltage for the most remote bus.
+Under this setting, the voltage profile from the weighted cost function results in an average deviation of $6.3\times 10^{-4}$ pu, which is significantly higher than that achieved with our proposed formulation ($2.3\times 10^{-4}$ pu).
+Moreover, the weighted formulation assigns higher costs to PV inverters located at the end of distribution feeders, leading to suboptimal and unfair utilization of available reactive power capacity.
+Consequently, it results in a total reactive power use of 460.4 kVarh, compared to 390.9 kVarh in our formulation.
+Reducing reactive power absorption can be beneficial for lowering system losses and extending PV inverter lifetime.
+In this simulation, the weighted cost function yields a loss of 65.4 kWh, which is higher than the 64.0 kWh loss achieved with our formulation.
+Finally, our formulation is flexible and can accommodate this alternative objective function if considered beneficial.
+IV-I Active power control
+Figure 16: Voltages for bus 95 under various objective functions with active power control.
+To further demonstrate the significance of the proposed formulation, we apply the algorithm for active power control to mitigate voltage limit violations in scenarios where reactive power capacity is insufficient.
+As shown in Fig.
+16, both formulations successfully enforce the voltage limit.
+However, the weighted formulation leads to 17.5% active power curtailment (1649.7 kWh generated out of a possible 2000.7 kWh), whereas the proposed formulation reduces curtailment to only 7.7% (1847.2 kWh generated).
+This highlights the practical advantage of the proposed approach.
 V Conclusion
@@
 The proposed distributed approach does not extend to congestion management while the centralized does so.
-The current presentation is based on radial systems but the approach can be readily extended to meshed systems using the approach in [3] to construct the sensitivity matrix $\mathbf{X}$.
 Extending the approach to unbalanced systems is an interesting future direction, where the focus is to construct a positive definite $\mathbf{X}$ with a sparse inverse matrix.
-Finally, the approach works for either active or reactive power control, but not joint control.
-This is because the observation in Section III-B does not hold anymore since the voltage difference is no longer the gradient of either active or reactive power injection, but represents their joint non-separable impact.
-However, one might deploy active and reactive power feedback controllers separately at different sampling time.
-Whether the system under two feedback controllers is stable is left for future investigations.
+Finally, the approach can be extended to joint active and reactive power control.